### v1

# Understanding Sampler Stochasticity in Training Diffusion Models for RLHF

Abstract
Reinforcement Learning from Human Feedback (RLHF) is increasingly used to fine-tune diffusion models, but a key challenge arises from the mismatch between stochastic samplers used during training and deterministic samplers used during inference. In practice, models are fine-tuned using stochastic SDE samplers to encourage exploration, while inference typically relies on deterministic ODE samplers for efficiency and stability. This discrepancy induces a reward gap, raising concerns about whether high-quality outputs can be expected during inference. In this paper, we theoretically characterize this reward gap and provide non-vacuous bounds for general diffusion models, along with sharper convergence rates for Variance Exploding (VE) and Variance Preserving (VP) Gaussian models. Methodologically, we adopt the generalized denoising diffusion implicit models (gDDIM) framework to support arbitrarily high levels of stochasticity, preserving data marginals throughout. Empirically, our findings through large-scale experiments on text-to-image models using denoising diffusion policy optimization (DDPO) and mixed group relative policy optimization (MixGRPO) validate that reward gaps consistently narrow over training, and ODE sampling quality improves when models are updated using higher-stochasticity SDE training.
1 Introduction
Diffusion models (e.g., Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), FLUX (Black Forest Labs, 2024)) have shown strong performance in text-to-image (T2I) tasks, and have also been extended beyond images to video (Ho et al., 2022) and audio (Liu et al., 2023). To meet downstream objectives such as aesthetics, safety, and alignment, it is essential to post-train with RLHF (Ouyang et al., 2022) for preference-driven improvements, often with a KL-regularization term to preserve performance on pretrained tasks (Schulman et al., 2017). Widely used RLHF algorithms include DDPO (Black et al., 2024) and GRPO (Shao et al., 2024) variants (FlowGRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), MixGRPO (Li et al., 2025)). DDPO directly optimizes human-preference rewards by casting the denoising process as a Markov Decision Process (MDP); GRPO variants use group-relative advantages. See (Winata et al., 2025) for a broader review of successful RLHF algorithms for generative models.
Despite strong progress in alignment, RLHF training often exhibits unstable trajectories, long inference times, and vulnerability to reward hacking (Skalse et al., 2022). Using multiple rewards can mitigate the latter (Lee et al., 2025), but we also need efficient, robust samplers to produce stable, high-quality fine-tuned models. A classical scheme is DDPM (Ho et al., 2020), a discretization of the score-based backward SDE (Risken, 1996; Song et al., 2021b), which preserves data marginals but is time-consuming. In contrast, the deterministic DDIM sampler (Song et al., 2021a) follows the probability-flow ODE, enabling marginal-preserving sampling with fewer steps. DDPO typically uses DDIM with constant stochasticity to generate training samples; MixGRPO mixes SDE and ODE steps, varying stochasticity across the denoising horizon. However, although stochasticity is valuable for generating diverse training data, fine-tuned models often use DDIM or higher-order ODE solvers (Lu et al., 2022; 2025) for fast, stable inference. This naturally raises the question:
Why can we guarantee good sampling quality when inference uses a different noise level than the one used during RLHF training?
SDE Refer to caption Refer to caption Refer to caption Refer to caption
ODE Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: ODE (below) image generation preserves prompt instructions with better quality on details compared to SDE (above) image generation under large stochasticity $(\eta=1.2)$.
Prompts (from left to right): “A vintage writing desk with an open journal and a flickering candle.”, “A macaque soaking in a steaming hot spring, surrounded by falling snow.”, “A wicker rocking chair on a wrap-around porch during golden hour.”, “A sleek sports car drifting on a mountain highway during golden hour.”
In this paper, we address it from three perspectives. Theoretically, we derive non-vacuous bounds on the reward gap between a generally SDE-fine-tuned model and its ODE-sampling counterpart using Gronwall’s inequality. Specifically, for VE and VP Gaussian models, the gap shrinks at sharp rates $O(1/T)$ and $O(e^{-T^{2}}/2)$, where $T$ is the denoising time horizon. These results justify the practice of ODE inference for SDE-fine-tuned models and provide insights into more randomized exploration beyond $\eta=1.0$, the noise level for continuous score-based backward dynamics. Methodologically, we adopt the gDDIM (Zhang et al., 2023) framework for arbitrary stochasticity levels. Rather than linear variance interpolations between DDIM and DDPM, our implementation preserves marginals for all $\eta\geq 0$ and, more importantly, supports $\eta>1$ in a principled way, enabling broader RLHF exploration beyond standard DDPM noise levels. Empirically, we evaluate large-scale T2I models and RLHF algorithms to quantify the reward gap across multiple reward functions. The results show that (i) reward gaps decrease as training quality improves, (ii) moderate-to-high training stochasticity (e.g., $\eta=1.2$) yields superior in- and out-of-domain performance, and (iii) ODE inference remains stable and often outperforms SDE inference under small denoising step budgets.
  • Theory: Sharp bounds for VE/VP models; non-vacuous bounds for general distributions (linear in stochasticity, bounded in time-steps).
  • Methodology: gDDIM-based noise scheduling that preserves marginals for all noise levels ($\eta\geq 0$) and supports high stochasticity ($\eta>1$) in a principled way.
  • Experiments: Comprehensive evaluations across rewards and RLHF algorithms, validating bounded reward gaps and improved ODE inference with higher training stochasticity.
Related Literature. Higher-order ODE solvers such as RX-DPM (Choi et al., 2025), DEIS (Zhang & Chen, 2022) approximate probability-flow. Recent analyses give broader convergence guarantees (Huang et al., 2025) and better representability (Chen et al., 2023) for score-based diffusion. In parallel, RL-based fine-tuning leverages stochastic sampling: DRaFT differentiates through noisy trajectories (Clark et al., 2023), Score-as-Action casts fine-tuning as stochastic control (Zhao et al., 2025), and Adjoint Matching enforces optimal memoryless noise schedules (Domingo-Enrich et al., 2024). Large-scale studies combine multiple rewards (Zhang et al., 2024); SEPO, D3PO, and ImageReFL boosts alignment (Zekri & Boullé, 2025; Yang et al., 2024; Sorokin et al., 2025). Finally, (Liang et al., 2025) gives discretization error bound and (Wu et al., 2024) discusses guidance (Ho & Salimans, 2021) for Gaussian Mixture priors.
Organization of the paper. The remainder of the paper is organized as follows. In Section 2, we provide background on diffusion models and the RLHF framework. The theory is developed in Sections 3 and 4. In Section 5, we conduct numerical experiments to justify the role of sampler stochasticity in fine-tuning large-scale T2I models. Concluding remarks are in Section 6.
2 Preliminaries
Figure 2: One-dimensional VP dynamics under different noise levels ($\eta$), using the same control function. Increasing $\eta$ introduces higher stochasticity in the trajectories while preserving the underlying dynamics.
2.1 Forward-Backward Diffusion Models
The continuous-time diffusion models (Song et al., 2021b) involve forward and backward procedures, and provide a framework that both unifies and extends earlier discrete-time models. The brief review below of this approach is based on (Tang & Zhao, 2025).
The goal of diffusion models is to generate new samples that resemble the target data distribution $p_{\mbox{data}}(\cdot)$. The forward process is given by an SDE:
$dX_{t}=f(t,X_{t})dt+g(t)dB_{t},\quad X_{0}\sim p_{\mbox{data}}(\cdot),$ (1)
where $\{B_{t}\}$ is the $d$-dimensional Brownian motion, and $f:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are given functions (model parameters). Some regularity conditions on $f(\cdot,\cdot)$ and $g(\cdot)$ are required to ensure equation 1 is well-defined, see (Stroock & Varadhan, 1979). Moreover, assume the forward process $\{X_{t}\}$ has a smooth density $p(t,x):=\mathbb{P}(X_{t}\in dx)/dx$. It is known (Anderson, 1982; Haussmann & Pardoux, 1986) that under time reversal, the process, denoted $\overline{X}_{t}=X_{T-t}$, is governed by the SDE:
$d\overline{X}_{t}=(-f(T-t,\overline{X}_{t})+g^{2}(T-t)\nabla\log p(T-t,\overline{X}_{t}))dt+g(T-t)dB_{t},\;\overline{X}_{0}\sim p(T,\cdot).$
There are two key points to note. First, as the goal is to generate the target distribution, which is unknown, the backward process is initiated with a prior $p_{\mbox{noise}}(\cdot)$, a proxy of $p(T,\cdot)$), instead of setting $\overline{X}_{0}\sim p(T,\cdot)$. The choice of $p_{\mbox{noise}}(\cdot)$ depends on each specific model, and varies case by case in applications. Similarly, the choice of the pair of functions $(f,g)$ also depends on applications, and notable examples include variance exploding (VE) and variance preserving (VP) SDEs (see Section 3).
Second, since the term $\nabla\log p(T-t,\overline{X}_{t})$ in the time-reversal process $\{\overline{X}_{t}\}$ is unknown, we need to find a way to estimate/approximate it. Specifically, calling $\nabla\log p(t,x)$ the “score function”, score matching refers to estimate the score function via a family of parameterized functions $\{s_{\theta}(t,x)\}_{\theta}$ (e.g., via neural nets). The most widely used approach is denoising score matching Vincent (2011):
$\min_{\theta}\mathbb{E}_{t\sim\mbox{Unif}\,[0,T],\,X_{0}\sim p_{data}(\cdot)}[\mathbb{E}_{p(t,\cdot|X_{0})}|s_{\theta}(t,X_{t})-\nabla\log p(t,X_{t}|X_{0})|^{2}].$
Replacing $\nabla\log p(t,x)$ by $s_{\theta}(t,x)$, the (backward) diffusion sampler is:
$dY_{t}=\left(-f(T-t,Y_{t})+g^{2}(T-t)s_{\theta}(T-t,Y_{t})\right)dt+g(T-t)dB_{t},\quad Y_{0}\sim p_{\mbox{noise}}(\cdot).$ (2)
For richer levels of trajectory variation while keeping terminal marginals, the sampler is set to:
$dY_{t}=\left(-f(T-t,Y_{t})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta}(T-t,Y_{t})\right)dt+\eta\,g(T-t)dB_{t},\quad Y_{0}\sim p_{\mbox{noise}}(\cdot),$ (3)
where $\eta\geq 0$ controls the sampler stochasticity along the trajectories. When $\eta=1$, equation 3 reduces to equation 2. When $\eta=0$, equation 3 becomes the probability flow ODE. We will use the processes specified by equation 3 as the pretrained models for fine-tuning. In particular, while prior studies focus on $\eta\in[0,1]$, we will explore the possible usage of $\eta>1$.
2.2 Discretization – Stochastic gDDIM
As shown in (Song et al., 2021b), the DDPM scheme is a discretized SDE that preserves variance, and the deterministic DDIM is a discretization of the probability flow ODE that preserves variance. For most diffusion models, the model parameters are:
$f(t,x)=\frac{1}{2}\frac{d\log\alpha_{t}}{dt}x\quad\mbox{and}\quad g(t)=\sqrt{-\frac{d\log\alpha_{t}}{dt}},$
where $\{\alpha_{t}\}_{t=0}^{T}$ is a decreasing sequence with $\alpha_{0}=1$ and $\alpha_{T}=0$. (Zhang et al., 2023) proposed the following discretization of equation 3 for all $\eta\geq 0$:
$x_{t-\Delta t}=\sqrt{\frac{\alpha_{t-\Delta t}}{\alpha_{t}}}x_{t}$ $+\left(\sqrt{\frac{\alpha_{t-\Delta t}}{\alpha_{t}}}(1-\alpha_{t})-\sqrt{(1-\alpha_{t-\Delta t}-\sigma_{t}^{2})(1-\alpha_{t})}\right)s_{\theta}(t,x_{t})$ (4)
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sigma_{t}(\eta)\mathcal{N}(0,I),\quad x_{T}\sim p_{\mbox{noise}}(\cdot).$
where $\{x_{t}\}_{t=0}^{T}$ is the discretized sequence of backward diffusion $Y_{t}$ and
$\sigma_{t}(\eta)=(1-\alpha_{t-\Delta t})\left(1-\left(\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}\right)^{\eta^{2}}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\right)^{\eta^{2}}\right).$ (5)
equation 4 is referred to as the generalized denoising diffusion implicit model (gDDIM), which agrees with the stochastic DDIM (Song et al., 2021a). The difference lies in the choice of $\sigma_{t}(\eta)$: (Song et al., 2021a) relied on heuristics to suggest $\sigma_{t}(\eta)\propto\eta$, whereas equation 5 proposed by (Zhang et al., 2023) guarantees the exact sampling if the score function is known/precise. When $\eta=0$, we have $\sigma_{t}(0)=0$, and equation 4 is just the deterministic DDIM. We use gDDIM scheme to conduct numerical experiments with controlled stochasticity, of which the results are detailed in Section 5.
2.3 Diffusion RLHF
RLHF was originally proposed for LLM alignment (Bai et al., 2022; Ouyang et al., 2022). Such a paradigm can also be made well-suited for fine-tuning diffusion models, particularly to enhance T2I generation aligned with human feedback. The first works in this direction are (Black et al., 2024; Fan et al., 2023; Lee et al., 2023), which formulated the denoising step as Markov decision processes (MDPs). (Gao et al., 2024; Zhao et al., 2024; 2025) developed an RL approach based on continuous-time models. Here we focus on the most widely used approach of Denoising Diffusion Policy Optimization (DDPO).
DDPO (Black et al., 2024): We formulate the denoising steps $\{x_{T},x_{T-1},\cdots,x_{0}\}$ as an MDP:
$s_{t}:=(\bm{c},t,x_{t}),$ $\pi(a_{t}|s_{t}):=p_{\theta}(x_{t-1}|x_{t},\bm{c}),\qquad$ $\mathbb{P}(s_{t+1}|s_{t},a_{t}):=\big(\delta_{\bm{c}},\delta_{t-1},\delta_{x_{t-1}}\big),$
$a_{t}:=x_{t-1},$ $\rho_{0}(s_{0}):=\big(p(\bm{c}),\delta_{T},\mathcal{N}(0,I)\big),\qquad$ $R(s_{t},a_{t}):=\begin{cases}r(x_{0},\bm{c})&\text{if }t=0,\\ 0&\text{otherwise},\end{cases}$
where $p_{\theta}$ is the probability from $x_{t}$ to $x_{t-1}$ given prompt $\bm{c}$ and $\theta$-parametrized score function $s_{\theta}$; $\delta_{\bullet}$ is the Dirac function that distributes the (non-zero) mass at $\bullet$ only; $R$ is the reward function that is equal to the denoised image quality $r(x_{0},c)$ for each trajectory. The objective of DDPO is to maximize:
$\mathcal{J}_{DDPO}(\theta):=\mathbb{E}_{\bm{c}\sim p(\bm{c}),x_{0}\sim p_{\theta}(x_{0}|\bm{c})}[r(x_{0},\bm{c})].$ (6)
The policy gradient of equation 6 is $\nabla_{\theta}\mathcal{J}_{DDPO}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},\bm{c})\,r(x_{0},\bm{c})\right]$ (Williams, 1992). It is common to parameterize $p_{\theta}(x_{t-1}|x_{t},\bm{c})$ as isotopic Gaussian, so $\nabla_{\theta}\mathcal{J}_{DDPO}$ can be easily computed by Monte Carlo (Mohamed et al., 2020).
GRPO (Shao et al., 2024): A recent breakthrough in the space of RLHF is the Group Relative Policy Optimization (GRPO) framework in the wake of Deepseek-R1 for reasoning. For T2I tasks, GRPO uses a group of sampled outputs (indexed by $i$ or $k$ below), along with the group normalization/average per prompt, to compute the advantages:
$A_{i}:=\frac{r(x_{0}^{i},\bm{c})-\frac{1}{G}\sum_{k=1}^{G}r(x_{0}^{k},\bm{c})}{\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)},$
where $G$ is the group size; $\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)$ denotes the standard deviation of the group rewards. The objective of GRPO is to maximize:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=1}^{T}\,\min\left(\rho_{t}^{i}(\theta)A_{i},\text{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (7)
where $\rho_{t}^{i}(\theta)=\frac{p_{\theta}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}{p_{\theta_{\mbox{old}}}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}$ is the likelihood ratio; $\{x_{t}^{i}\}$ is the $i$-th sample trajectory in the group; and the expectation is with respect to $\bm{c}\sim p(\bm{c})$ and $\{x^{i}_{t}\}\sim\pi_{\theta_{\mbox{old}}}(\cdot|\bm{c})$. Variants of the GRPO algorithm (Xue et al., 2025; Li et al., 2025) have been proposed for T2I tasks.
3 Theory on the reward gap
As mentioned in the last section, the forward-backward diffusion approach to generative AI usually makes use of stochasticity for exploration (the forward step), while relying on deterministic samplers for inference (the backward step); the latter is obviously motivated by computational efficiency. Yet, there’s no study to examine/analyze the role of stochasticity; namely, whether there’s a right or “best” level of stochasticity to use, in both steps; and in particular, whether it’d be worthwhile to use stochasticity to the backward step, at the price of reduced efficiency.
Denote by $\{Y^{REF}_{t}\}$ the “reference” model specified in equation 3), with a specific choice of the parameter $\eta$. Denote by $\{Y^{SDE}_{t}\}:=\{Y^{\theta}_{t}\}$ the fine-tuned model, score functions parameterized by $\theta$.
To carry out this analysis, consider the following entropy-regularized reward objective (e.g., (Uehara et al., 2024; Tang, 2024)):
$F_{\eta}(\theta)=\mathbb{E}[r(Y^{\theta}_{T})]-\beta\,\text{KL}(Y_{T}^{\theta}||Y_{T}^{REF}),$ (8)
where $r(\cdot)$ is a reward function, and $\beta>0$ controls the level of exploration relative to the pretrained model. $Y^{SDE}_{t}$ evolves under a same $\eta$ as $Y^{REF}_{t}$, and the problem is to solve $\theta^{*}_{\eta}:=\arg\max_{\theta}F_{\eta}(\theta)$.
$\bullet$ Inference: Denote by $\{Y_{t}^{ODE}\}$ the marginal-preserved (deterministic) ODE sampler by letting $\eta\downarrow 0$ in $\{Y^{SDE}_{t}\}$.
Also let $\mathcal{J}_{\bullet}:=\mathbb{E}[r(Y_{T}^{\bullet})]$ be the evaluation metric, with $\bullet\in\{REF,SDE,ODE\}$. The goal is to understand how much improvement gained from fine-tuning, and the reward gap induced by the (deterministic) ODE sampler. This motivates the following definitions:
Definition 3.1.
Define
  1. 1.
    Improvement of fine-tuning as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{REF}_{t}\}$, $\mathcal{}I_{\eta}:=\mathcal{J}_{ODE}-\mathcal{J}_{REF}\geq 0$.
  2. 2.
    Reward gap as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{SDE}_{t}\}$, $\Delta_{\eta}:=|\mathcal{J}_{ODE}-\mathcal{J}_{SDE}|$.
In what follows, we study $I_{\eta}$ and $\Delta_{\eta}$ for the VE/VP models with a Gaussian or mixture Gaussian target distribution, where the score function has a closed-form expression. A non-vacuous bound for the general case will be given in the next section.
Remark.
$\mathcal{J}_{ODE}\geq\mathcal{J}_{REF}$ for all models, following from the optimality of $\theta^{*}_{\eta}$. $\mathcal{J}_{ODE}>\mathcal{J}_{SDE}$ for VE/VP models and DDPO; $\mathcal{J}_{SDE}>\mathcal{J}_{ODE}$ for MixGRPO.
3.1 VE with a Gaussian target
We first consider the one-dimensional VE model:
$dX_{t}=\sqrt{2t}\,dB_{t},\quad\mbox{with }\,p_{\mbox{data}}(\cdot)=\mathcal{N}(0,1).$ (9)
Since $X_{t}\sim\mathcal{N}(0,t^{2}+1)$, the (exact) score function is $\nabla\log p(t,x)=-\frac{x}{t^{2}+1}$. So the “pretrained” model is:
$dY_{t}^{REF}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t}.$
Next we set the reward function $r(x)=-(x-1)^{2}$, so the goal of fine-tuning is to drive the sample towards the mode $1$. We also specify the fine-tuned SDE and ODE by
$dY_{t}^{SDE}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}(Y_{t}^{SDE}+\theta_{\eta}^{*})dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{ODE}=-\frac{T-t}{1+(T-t)^{2}}(Y_{t}^{ODE}+\theta_{\eta}^{*})dt,$
with $\theta_{\eta}^{*}$ maximizing the entropy-regularized reward (equation 8). The following theorem gives bounds for $I_{\eta}$ and $\Delta_{\eta}$ under VE, and the proof is deferred to Appendix C.1.
Theorem 3.1.
Consider the Variance Exploding model (equation 9), with the reward function
$r(x)=-(x-1)^{2}.$ (10)
For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\right)\right]^{-1}$. Moreover,
$0\leq\Delta_{\eta}\leq\frac{1}{2T}+o\left(\frac{1}{T}\right)\quad\mbox{and}\quad\mathcal{I}_{\eta}\geq 1-\frac{1}{2T}+o\left(\frac{1}{T}\right).$ (11)
3.2 VP with a Gaussian target
Now we consider the one-dimensional VP model:
$dX_{t}=-tX_{t}dt+\sqrt{2t}\,dB_{t},\quad\mbox{with }\,p_{\mbox{data}}(\cdot)=\mathcal{N}(0,1).$ (12)
Under the same setup as in the VE case, we have:
$dY_{t}^{REF}=-\eta^{2}(T-t)Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{SDE}=-\eta^{2}(T-t)Y_{t}^{SDE}\,dt-(1+\eta^{2})(T-t)\theta^{*}_{\eta}(t)\,dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{ODE}=-(T-t)\theta_{\eta}^{*}(t)dt,$
where a time-dependent control $\theta_{\eta}(t):=\theta_{\eta}e^{-\frac{(T-t)^{2}}{2}}$ is used for fine-tuning. The following theorem gives bounds for $I_{\eta}$ and $\Delta_{\eta}$ under VP, and the proof is deferred to Appendix C.2.
Theorem 3.2.
Consider the Variance Preserving model (equation 12), with the reward function $r(x)=-(x-1)^{2}$. For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-e^{-\frac{(1+\eta^{2})T^{2}}{2}}\right)\right]^{-1}$. Moreover,
$0\leq\Delta_{\eta}\leq\frac{e^{-T^{2}}}{2}+o\left(e^{-T^{2}}\right)\quad\mbox{and}\quad\mathcal{I}_{\eta}\geq 1-\frac{e^{-T^{2}}}{2}+o\left(e^{-T^{2}}\right).$ (13)
3.3 VE/VP with a mixture Gaussian target
The previous results can be extended to multidimensional setting, with a mixture Gaussian target distribution. Recall that the probability density of a mixture Gaussian has the form:
$\sum_{i=1}^{k}\frac{\alpha_{i}}{(2\pi)^{d/2}(\det{\bm{\Sigma}_{i}})^{1/2}}\cdot\exp\left(-\frac{1}{2}(x-\bm{\mu}_{i})^{T}{\bm{\Sigma}_{i}}(x-\bm{\mu}_{i})\right),$
where $\alpha_{i}$ is the weight of the $i$-th Gaussian component. The following corollary bounds the reward gap for a mixture Gaussian target distribution, and the proof is deferred to Appendix C.3.
Corollary 3.1.
[…]
$\Delta_{\eta}\leq\left\{\begin{array}[]{ll}(2T)^{-1}&\text{for VE},\\ e^{-T^{2}}/2&\text{for VP}.\end{array}\right$. (14)
For all these examples, the reward gap $\Delta_{\eta}\to 0$ (independent of $\eta$), as $T\to\infty$. Such a phenomenon will also be observed in fine-tuning T2I models with more complex rewards.
4 Non-vacuous bound on the reward gap
In Section 3, we studied the reward gap for the VE and VP models with a Gaussian/mixture Gaussian target distribution. The analysis relies on the explicit computation of the score function, which is not available in general. Here our goal is to provide a bound on $W_{2}(Y_{T}^{SDE},Y_{T}^{ODE})$, where
$dY_{t}^{SDE}=\left(-f(T-t,Y_{t}^{SDE})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{SDE})\right)dt+\eta\,g(T-t)dB_{t},$
$dY_{t}^{ODE}=\left(-f(T-t,Y_{t}^{ODE})+\frac{1}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{ODE})\right)dt.$
Here we fine-tune directly on the score matching function $\{s_{\theta}(t,x)\}_{\theta}$, instead of adding a control as in Section 3. We need the following assumptions.
Assumption 4.1.
The following conditions hold:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
The conditions 1 (dissipativity of $f$) and 2 (Lipschitz condition on $s_{\theta}$) are standard in stability analysis. The condition 3 ensures process $\{Y^{SDE}_{t}\}$ a contractive drift, which holds if $f$ is dissipative, and the fine-tuned distribution satisfies some strongly log-concave condition, see e.g., (Gao et al., 2025; Tang & Zhao, 2024). The following theorem gives a non‑vacuous bound on the reward gap, see proof in Appendix D.
$\eta$ ImageReward PickScore HPS_v2 Aesthetic
Base
0.320 20.69 0.253 5.375
ImageReward
1.0 0.837 (0.918) 20.89  (20.94) 0.268 (0.271) 5.638  (5.720)
1.2 0.915 (1.031) 20.95  (20.99) 0.268 (0.289) 5.697 (5.803)
1.5 0.771 (0.907) 20.90 (20.95) 0.266 (0.269) 5.605 (5.735)
PickScore 1.0 0.587 (0.729) 21.11 (21.28) 0.265 (0.267) 5.560 (5.678)
1.2 0.594 (0.692) 21.17 (21.33) 0.264 (0.264) 5.568 (5.692)
1.5 0.566 (0.701) 21.10 (21.36) 0.264 (0.262) 5.599 (5.769)
Table 1: Performance for ODE (SDE) samplers under DDPO fine-tuning. Bold numbers indicate the highest evaluations among all stochasticity, demonstrating $\eta=1.2$ in general performs the best.
Theorem 4.1.
[…]
5 Numerical experiments
In this section, we conduct numerical experiments for fine-tuning large-scale T2I models by leveraging sampler stochasticity, and examine the reward gap. Two classes of RL algorithms are considered: DDPO (Black et al., 2024) and MixGRPO (Li et al., 2025). The preference rewards that we use for fine-tuning DDPO include the LAION aesthetic (Schuhmann et al., 2022), HPS_v2.1 (Wu et al., 2023), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2023). The rewards for fine-tuning MixGRPO include HPS_v2.1, PickScore, ImageReward, and Unified Reward. (Wang et al., 2025) To study the reward gap, we use a high stochasticity $\eta$ under gDDIM scheme to generate training samples, and compare them with the ODE samples generated using the same weights in each iteration. The stochasticity scale $\eta$ controls the backward dynamics according to equation 4 and equation 5 in Section 2.2. Choices of hypermaraters are detailed in Appendix E.
5.1 DDPO
We use Stable Diffusion v1.5 (Rombach et al., 2022) as the base model, and fine-tune it with DDPO. During training, we adopt ImageReward and PickScore as preference rewards, while Aesthetic and HPS_v2 are included as additional evaluation metrics. Figure 1 (and Appendix G) provide representative generations from the fine-tuned models. Our main observations are summarized as follows:
$\bullet$ Decreasing Reward Gap with Quality Improvement: we conduct fine-tuning experiments under PickScore with $\eta=1.2$ and calculate SDE–ODE reward (contrary to Section 3, here we subtract ODE reward value from SDE reward for non-negativity) differences under multiple preference functions every 200 steps until reward collapse. As shown in Table 2, the gap decreases as image quality improves for both samplers.
$\bullet$ High Stochasticity Benefits Moderate Time Steps: we compare fine-tuning under ImageReward and PickScore at $\eta\in\{1.0,1.2,1.5\}$ for 200 steps. As shown in Table 1, $\eta=1.2$ under ImageReward achieves the best in-domain and out-of-domain performance, while PickScore’s performances depend on evaluation metrics.
$\bullet$ Richer Prompt Contents Reduce Reward Gap: we compare performances with animal versus more comprehensive prompts (Figure 1, Appendix G) under ImageReward with $\eta=1.2$. As shown in Table 3, complex prompts generate higher post-tuning rewards and higher in-group variance in post-training. Moreover, their richer instructions reduce the SDE–ODE reward gap.
Refer to caption Refer to caption
Figure 3: Evolution of reward gap during DDPO training under PickScore fine-tuning with stochasticity $\eta\in\{1.0,1.2,1.5\}$. The gap for multiple results decreases steadily as training progresses, indicating that image quality improves for both samplers and that ODE inference remains competitive.
$T=\bm{0}$ 200 400 600 800 1000 1200 1400 (1476)
ImgRwrd Gap 0.160 0.119 0.102 0.028 0.057 0.027 0.006 0.020 (0.564)
HPSv2 Gap 0.0047 0.0032 0.0032 0.0018 0.0027 0.0015 -0.0028 0.0011 (0.0308)
Aesthetic Gap 0.162 0.113 0.078 0.077 0.072 0.017 0.030 -0.010 (0.685)
PickScore Gap 0.115 0.146 0.167 0.118 0.178 0.106 0.129 0.090 (0.907)
SDE Reward 20.823 21.310 21.647 21.901 22.068 22.265 22.306 22.417 (18.920)
ODE Reward 20.730 21.165 21.500 21.783 21.883 22.172 22.195 22.318 (17.044)
Table 2: PickScore training until reward collapses. Smallest reward gaps and best sampler performances locate at the large training steps $T=1200,1400$.
Animal Prompts Comprehensive Prompts
$T=\bm{0}$ 100 200  $T=\bm{0}$ 100 200
Mean 0.382 (0.539) 0.668 (0.805) 0.915 (1.031)    0.347 (0.470) 0.763 (0.872) 1.086 (1.174)
Std 0.814 (0.799) 0.775 (0.721) 0.709 (0.652)    1.030 (0.997) 0.909 (0.868) 0.795 (0.727)
Gap 0.146 0.138 0.116  0.115 0.110 0.106
Table 3: Performance Comparison between prompts of different complexity. More complicated prompts yields faster fine-tuning improvements, larger in-group variances, and smaller SDE-ODE reward gaps.
5.2 MixGRPO
We use FLUX.1 (Black Forest Labs, 2024) as the base model, and fine-tune it with MixGRPO, which is a sliding-window sampler that alternates between ODE and SDE schemes for 25 training steps in total. Training is carried out with multiple rewards combined using equal weights, while evaluation is reported on ImageReward and HPSClip.
$\bullet$ Bounded Reward Gap Under High Stochasticity: With $\eta=1.2$, ODE sampling in MixGRPO consistently outperforms the mixed SDE–ODE scheme, in contrast to the results from DDPO. As shown in the top-left and bottom-left panels of Figure 4, the reward gap steadily diminishes during training and converges to zero.
$\bullet$ Quality Improvement: The middle and right panels of Figure 4 further shows that the ODE sampler performs consistently better on training prompts. For example, in the left pair of Figure 5, the generation with $\eta=1.2$ correctly aligns with the “trapped inside” prompt instruction, whereas the generation with $\eta=1.0$ fails to do so.
Figure 4: Bounded reward gap (left column) and performance improvement (middle, right column) for MixGRPO
Noise Level = 1.0 Refer to caption Refer to caption Refer to caption
Noise Level = 1.2 Refer to caption Refer to caption Refer to caption
Figure 5: Comparison of ODE image generation by FLUX with MixGRPO fine-tuning, stochasticity $\eta=1.2$ (below) and $\eta=1.0$ (above). Higher stochasticity shows better alignments to details.
Prompts (from left to right): “A steampunk pocketwatch owl is trapped inside a glass jar buried in sand, surrounded by an hourglass and swirling mist.”, “An androgynous glam rocker poses outside CBGB in the style of Phil Hale.”, “A digital painting by Loish featuring a rush of half-body, cyberpunk androids and cyborgs adorned with intricate jewelry and colorful holographic dreads.”
6 Conclusion and Further Directions
Sampling stochasticity in diffusion RLHF plays an important role yet remains underexamined. Here we present an explicit solution for generalized VE/VP diffusion models, together with extensive experiments, demonstrating that “high stochasticity in training samples, no stochasticity in generation” is both theoretically sound and practically advantageous. Our results open a broader space for tuning stochasticity hyperparameters, enabling more robust and diversified diffusion post-training. Future work will quantify the “reward gap” in video and multimodal generation and develop clearer theoretical insights into stochasticity as a drifting force toward the target distribution.
Appendix A Lemmas on Linear Dynamics with Gaussian priors
Lemma A.1.
A stochastic process $\{Z_{t}\}_{t=0}^{T}$ with first order linear dynamic and initial Gaussian distribution
$\left\{\begin{array}[]{ll}dZ_{t}=f(t)Z_{t}dt+g(t)dt+h(t)dB_{t}&t\in[0,T]\\[3.00003pt] Z_{0}\sim\mathcal{N}(0,1)\end{array}\right$.
is distributed following
$Z_{t}\sim\mathcal{N}\Big(e^{F(t)}\int_{0}^{t}e^{-F(s)}g(s)\,ds,\;e^{2F(t)}\int_{0}^{t}e^{-2F(t)}h^{2}(s)\,ds\Big),$
in which $F(t)$ is the integrating factor satisfying $F(t)=\int_{0}^{t}f(s)ds.$
Lemma A.2.
A stochastic process $\{Z_{t}\}_{t=0}^{T}$ with first order linear dynamic and initial Gaussian distribution
$\left\{\begin{array}[]{ll}dZ_{t}=f(t)Z_{t}dt+g(t)dB_{t}&t\in[0,T]\\[3.00003pt] Z_{0}\sim\mathcal{N}(\mu_{Z},\sigma_{Z}^{2})\end{array}\right$.
is distributed following
$Z_{t}$ $\sim Z_{0}\cdot e^{F(t)}+\mathcal{N}\Big(0,\int_{0}^{t}e^{-2F(s)}g^{2}(s)ds\Big)\cdot e^{F(t)}$
$\sim\mathcal{N}\Big(\mu_{Z}\cdot e^{F(t)},\Big[\sigma_{Z}^{2}+\int_{0}^{t}e^{-2F(s)}g^{2}(s)ds\Big]\cdot e^{2F(t)}\Big),$
in which $F(t)$ is the integrating factor satisfying $F(t)=\int_{0}^{t}f(s)ds.$
Lemma A.3.
A parametrized family of stochastic processes $\{Z_{t}^{\theta}\}_{t=0}^{T}$ with initial Gaussian distribution
$\left\{\begin{array}[]{ll}dZ_{t}^{\theta}=f(t)\cdot(Z_{t}^{\theta}+\theta(t))dt+g(t)dB_{t}&t\in[0,T]\\[3.00003pt] Z_{0}\sim\mathcal{N}(\mu_{Z},\sigma_{Z}^{2})\end{array}\right$.
is distributed following
$Z_{t}$ $\sim Z_{0}\cdot e^{F(t)}+\mathcal{N}\Big(\int_{0}^{t}e^{-F(s)}f(s)\theta(s)ds,\int_{0}^{t}e^{-2F(s)}g^{2}(s)ds\Big)\cdot e^{F(t)}$
$\sim\mathcal{N}\Big(\mu_{Z}\cdot e^{F(t)}+\int_{0}^{t}e^{-F(s)}f(s)\theta(s)ds,\Big[\sigma_{Z}^{2}+\int_{0}^{t}e^{-2F(s)}g^{2}(s)ds\Big]\cdot e^{2F(t)}\Big),$
in which $F(t)$ is the integrating factor satisfying $F(t)=\int_{0}^{t}f(s)ds.$
Appendix B Useful Propositions
Proposition B.1.
The terminal distribution of Variance Exploding parametrized backward dynamics $\{Y_{t}^{\theta}\}_{t=0}^{T}$ is always Gaussian following the law
$\mathcal{N}(\mu_{Y_{T}^{\theta}},\sigma^{2}_{Y_{T}}):=\mathcal{N}\Big(\theta\cdot(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-\theta,1-(1+T^{2})^{-(1+\eta^{2})}\Big).$
Therefore, reward function (10) takes the form of
$\mathcal{J}_{\eta}(\theta)=-\Big(\sigma_{Y_{T}}^{2}+(1-\mu_{Y_{T}^{\theta}})^{2}\Big).$
Proof.
By comparing the coefficients with Lemma A.1, we have
$\left\{\begin{array}[]{ll}f_{\eta}(t)=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}\\[3.00003pt] g_{\eta}(t)=\eta\sqrt{2(T-t)}\\[3.00003pt] \sigma_{Z}^{2}=T^{2}.\end{array}\right$.
Therefore, we first examine the exponential of integrating factor,
$e^{F_{\eta}(t)}$ $=\exp\left(\int_{0}^{t}f_{\eta}(s)ds\right)$
$=\exp\left(\int_{0}^{t}-\frac{(1+\eta^{2})(T-s)}{1+(T-s)^{2}}ds\right)$
$=\exp\left(\int_{1+T^{2}}^{1+(T-t)^{2}}\frac{(1+\eta^{2})d(1+(T-s)^{2})}{2(1+(T-s)^{2})}\right)$
$=\Big(\frac{1+T^{2}}{1+(T-t)^{2}}\Big)^{-\frac{1+\eta^{2}}{2}}.$
At terminal time $T$, the cumulative factor is
$e^{F_{\eta}(T)}=\exp\left(\int_{0}^{T}f_{\eta}(s)ds\right)=\big(1+T^{2}\big)^{-\frac{1+\eta^{2}}{2}}.$
Also, the cumulative Gaussian variance generated from the backward process is,
$\int_{0}^{t}e^{-2F_{\eta}(s)}g^{2}_{\eta}(s)ds$ $=\int_{0}^{t}\Big(\frac{1+T^{2}}{1+(T-s)^{2}}\Big)^{(1+\eta^{2})}\cdot(2\eta^{2}(T-s))ds$
$=\int_{1+T^{2}}^{1+(T-t)^{2}}(-\eta^{2})\Big(\frac{1+T^{2}}{1+(T-s)^{2}}\Big)^{(1+\eta^{2})}d(1+(T-s)^{2})$
$=(1+T^{2})^{(1+\eta^{2})}\int_{1+T^{2}}^{1+(T-t)^{2}}(-\eta^{2})(1+(T-s)^{2})^{-(1+\eta^{2})}\,d(1+(T-s)^{2})$
$=(1+T^{2})^{(1+\eta^{2})}\cdot\Big((1+(T-t)^{2})^{-\eta^{2}}-(1+T^{2})^{-\eta^{2}}\Big).$
At terminal time $T$, the variance from the process is
$\int_{0}^{T}e^{-2F_{\eta}(s)}g^{2}_{\eta}(s)ds$ $=(1+T^{2})^{(1+\eta^{2})}\cdot\Big(1-(1+T^{2})^{-\eta^{2}}\Big)$
$=(1+T^{2})^{(1+\eta^{2})}-(1+T^{2}).$
Together with the initial Gaussian variance, the terminal distribution remains a zero-mean Gaussian:
$Y_{T}$ $\sim\mathcal{N}\Big(0,\big[(1+T^{2})^{(1+\eta^{2})}-(1+T^{2}-\sigma_{Z})\big]\cdot(1+T^{2})^{-(1+\eta^{2})}\Big)$
$\sim\mathcal{N}\Big(0,\big[(1+T^{2})^{(1+\eta^{2})}-1\big]\cdot(1+T^{2})^{-(1+\eta^{2})}\Big)$
$\sim\mathcal{N}\Big(0,1-(1+T^{2})^{-(1+\eta^{2})}\Big).$
Now we consider the terminal distribution for the parametrized process $Y_{T}^{\theta}$. To reduce the problem to a dynamic with linear drift term, we define
$Z_{t}^{\theta}=Y_{t}^{\theta}+\theta.$
so that
$\left\{\begin{array}[]{ll}dZ_{t}=f_{\eta}(t)Z_{t}dt+g_{\eta}(t)dB_{t}&t\in[0,T]\\[3.00003pt] Z_{0}\sim\mathcal{N}(\theta,T^{2})\end{array}\right$.
Observe that both the dynamic variance and the initial distribution variance for $\{Z^{\theta}_{t}\}$ and $\{Y_{t}\}$ are the same, we may directly apply Lemma A.1 to obtain
$Z_{T}^{\theta}\sim\mathcal{N}\big(\theta\cdot(1+T^{2})^{-\frac{1+\eta^{2}}{2}},1-(1+T^{2})^{-(1+\eta^{2})}\big)$
Therefore,
$Y_{T}^{\theta}\sim\mathcal{N}\big(\theta\cdot(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-\theta,1-(1+T^{2})^{-(1+\eta^{2})}\big),$
and thus
$\left\{\begin{array}[]{ll}\mu_{Y_{T}^{\theta}}:=\theta\cdot(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-\theta\\[3.00003pt] \sigma^{2}_{Y_{T}}:=1-(1+T^{2})^{-(1+\eta^{2})}\end{array}\right$. (16)
In addition, we can now give a closed form representation of our reward,
$\mathcal{J}_{\eta}(\theta)$ $=\mathbb{E}[(Y_{T}^{\theta}-1)^{2}]$
$=\mathbb{E}[Y_{T}^{2}]-2\mathbb{E}[Y_{T}]+1$
$=\sigma_{Y_{T}}^{2}+\mu_{Y_{T}^{\theta}}^{2}-2\mu_{Y_{T}^{\theta}}+1$
$=\sigma_{Y_{T}}^{2}+\big(1-\mu_{Y_{T}^{\theta}}\big)^{2},$
which yields to the desired expression. ∎
Proposition B.2.
Given $\beta,\eta$, $T$, the unique maximizer to the entropy regularized target (8) is:
$\theta^{*}_{\eta}=-\Big((1+\frac{\beta}{2})\cdot\big[1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\big]\Big)^{-1}.$
Proof.
A classical distance result on two Gaussian distributions (Hershey & Olsen, 2007) states:
Lemma B.1.
The KL divergence for two Gaussian distributions $P\sim\mathcal{N}(\mu_{1},\sigma_{1})$ and $Q\sim\mathcal{N}(\mu_{2},\sigma_{2})$,
$\operatorname{KL}(P||Q)=\log\Big(\frac{\sigma_{2}}{\sigma_{1}}\Big)+\frac{\sigma_{1}^{2}+(\mu_{1}-\mu_{2})^{2}}{2\sigma_{2}^{2}}-\frac{1}{2}.$
In our model, $P\sim Y_{T}^{\theta}$ and $Q\sim Y_{T}^{\odot}$, so $\mu_{1}=\mu_{{Y_{T}^{\theta}}}$, $\sigma_{1}=\sigma_{Y_{T}}$, $\mu_{2}=0$, $\sigma_{2}=1$.
$\operatorname{KL}(Y_{T}^{\theta}||Y_{T}^{\odot})$ $=\log(\frac{1}{\sigma_{Y_{T}}})+\frac{\sigma_{Y_{T}}^{2}+\mu_{{Y_{T}^{\theta}}}^{2}-1}{2}$
$=-\log(\sigma_{Y_{T}})+\frac{\sigma_{Y_{T}}^{2}+\mu_{{Y_{T}^{\theta}}}^{2}-1}{2}.$
Therefore,
$-F_{\eta}(\theta)$ $=\sigma_{Y_{T}}^{2}+\big(\mu_{Y_{T}^{\theta}}-1\big)^{2}+\big(-\beta\log(\sigma_{Y_{T}})+\frac{\beta}{2}\cdot(\sigma_{Y_{T}}^{2}+\mu_{Y_{T}^{\theta}}^{2}-1)\big)$
$=\big(\sigma_{Y_{T}}^{2}-\beta\log(\sigma_{Y_{T}})+\frac{\beta}{2}\sigma_{Y_{T}}^{2}-\frac{\beta}{2}\big)+\big(\mu_{Y_{T}^{\theta}}-1\big)^{2}+\frac{\beta}{2}\mu_{Y_{T}^{\theta}}^{2}$
$=\big(\sigma_{Y_{T}}^{2}-\beta\log(\sigma_{Y_{T}})+\frac{\beta}{2}\sigma_{Y_{T}}^{2}-\frac{\beta}{2}+1\big)+\big(1+\frac{\beta}{2}\big)\mu_{Y_{T}^{\theta}}^{2}-2\mu_{Y_{T}^{\theta}}.$
So it suffices to minimize a quadratic function w.r.t $\mu_{Y_{T}^{\theta}}$, of which we know
$\mu_{Y_{T}^{\theta^{*}}}=-\frac{-2}{2(1+\frac{\beta}{2})}=(1+\frac{\beta}{2})^{-1};$
and thus
$\theta_{\eta}^{*}\cdot\big[(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-1\big]=\mu_{Y_{T}^{\theta^{*}}}=(1+\frac{\beta}{2})^{-1}.$
This gives us the unique maximizer
$\theta^{*}_{\eta}=\Big((1+\frac{\beta}{2})\cdot\big[(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-1\big]\Big)^{-1}$
as desired. ∎
Proposition B.3.
The terminal distribution of Variance Preserving parametrized backward dynamics $\{Y_{t}^{\theta}\}_{t=0}^{T}$ is always Gaussian following the law
$\mathcal{N}(\mu_{Y_{T}^{\theta}},1):=\mathcal{N}\Big(\theta\cdot e^{-\frac{(1+\eta^{2})\cdot T^{2}}{2}}-\theta,1\Big).$
And reward function (10) takes the form of $\mathcal{J}_{\eta}(\theta)=-\Big(1+(1-\mu_{Y_{T}^{\theta}})^{2}\Big).$
Proof.
By comparing with the coefficients in Lemma A.2, we have
$\left\{\begin{array}[]{ll}f_{\eta}(t)=-\eta^{2}(T-t)\\[3.00003pt] g_{\eta}(t)=-(1+\eta^{2})(T-t)e^{-\frac{(T-t)^{2}}{2}}\theta_{\eta}^{*}\\[3.00003pt] h_{\eta}(t)=\eta\sqrt{2(T-t)}\end{array}\right$.
Therefore, we first examine the exponential of integrating factor,
$e^{F_{\eta}(t)}$ $=\exp\left(\int_{0}^{t}f_{\eta}(s)ds\right)$
$=\exp\left(-\eta^{2}\int_{T-t}^{T}s\,ds\right)$
$=\exp\left(-\frac{\eta^{2}}{2}\cdot\left(T^{2}-(T-t)^{2}\right)\right).$
At terminal time $T$, the cumulative factor is
$e^{F_{\eta}(T)}=\exp\left(-\frac{\eta^{2}}{2}\cdot\left(T^{2}-(T-T)^{2}\right)\right)=e^{-\frac{\eta^{2}T^{2}}{2}}.$
Now we are able to compute
$\int_{0}^{t}e^{-F_{\eta}(s)}g(s)\,ds$ $=\int_{0}^{t}e^{\frac{\eta^{2}}{2}\left(T^{2}-(T-s)^{2}\right)}\cdot\left(-(1+\eta^{2})(T-s)e^{-\frac{(T-s)^{2}}{2}}\theta_{\eta}^{*}\right)ds$
$=(1+\eta^{2})\cdot\theta_{\eta}^{*}\cdot\int_{T-t}^{T}e^{\frac{\eta^{2}}{2}\left(T^{2}-s^{2}\right)}\cdot\left(se^{-\frac{s^{2}}{2}}\right)ds$
$=(1+\eta^{2})\cdot\theta_{\eta}^{*}\cdot\int_{T-t}^{T}e^{\frac{\eta^{2}T^{2}}{2}}\cdot\left(se^{-\frac{(1+\eta^{2})s^{2}}{2}}\right)ds$
$=(1+\eta^{2})\cdot\theta_{\eta}^{*}\cdot e^{\frac{\eta^{2}T^{2}}{2}}\cdot\left[\frac{1}{1+\eta^{2}}\cdot e^{-\frac{(1+\eta^{2})s^{2}}{2}}\right]_{s=T-t}^{s=T}$
$=\theta_{\eta}^{*}\cdot e^{\frac{\eta^{2}T^{2}}{2}}\cdot\left(e^{-\frac{(1+\eta^{2})T^{2}}{2}}-e^{-\frac{(1+\eta^{2})(T-t)^{2}}{2}}\right)$
Therefore,
$\mu_{Y_{T}^{\theta}}=e^{-\frac{\eta^{2}T^{2}}{2}}\cdot\theta_{\eta}^{*}\cdot e^{\frac{\eta^{2}T^{2}}{2}}\cdot(e^{-\frac{(1+\eta^{2})T^{2}}{2}}-e^{0})=\theta_{\eta}^{*}\cdot(e^{-\frac{(1+\eta^{2})T^{2}}{2}}-1)$
To the variance preserving property, it suffices to show
$\frac{1}{e^{2F_{\eta}(t)}}=\int_{0}^{t}e^{-2F_{\eta}(t)}h_{\eta}^{2}(s)\,ds.$
In fact,
$\frac{d}{dt}e^{-2F_{\eta}(t)}=e^{-2F_{\eta}(t)}\cdot\frac{d(-\eta^{2}(T-t)^{2})}{dt}=e^{-2F_{\eta}(t)}\cdot\left(\eta^{2}\cdot 2(T-t)\right)=e^{-2F_{\eta}(t)}\cdot h_{\eta}^{2}(t).$
Proposition B.4.
Given $\beta,\eta$, $T$, the unique maximizer to the entropy regularized target (8) is:
$\theta^{*}_{\eta}=-\Big((1+\frac{\beta}{2})\cdot\big[1-e^{-\frac{(1+\eta^{2})\cdot T^{2}}{2}}\big]\Big)^{-1}.$
Proof.
Since $\sigma\equiv 1$, according to Lemma B.1, the maximum reward is attained at
$\mu_{Y_{T}^{\theta^{*}}}=-\frac{-2}{2(1+\frac{\beta}{2})}=(1+\frac{\beta}{2})^{-1}.$
By Proposition B.1,
$\theta_{\eta}^{*}\cdot\big[e^{-\frac{(1+\eta^{2})\cdot T^{2}}{2}}-1\big]=\mu_{Y_{T}^{\theta^{*}}}=(1+\frac{\beta}{2})^{-1}.$
This gives us the unique maximizer
$\theta^{*}_{\eta}=\Big((1+\frac{\beta}{2})\cdot\big[e^{-\frac{(1+\eta^{2})\cdot T^{2}}{2}}-1\big]\Big)^{-1}.$
Appendix C Proofs for Section 3
C.1 Proof of Theorem 3.1
We first consider the $Y^{ODE}_{T}$ process as discussed in Section 3.1. Since $\eta=0$, by Proposition B.2,
$\mu_{Y_{T}^{ODE}}=\theta_{\eta}^{*}\cdot(1+T^{2})^{-\frac{1}{2}}-\theta_{\eta}^{*}=(1+\frac{\beta}{2})^{-1}\cdot\frac{(1+T^{2})^{-\frac{1}{2}}-1}{(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-1},$
and
$\sigma^{2}_{Y_{T}^{ODE}}=\sigma^{2}_{Y_{T}^{0}}=1-(1+T^{2})^{-1}.$
Moreover, the quadratic reward $\mathcal{J}_{ODE}$ for $Y_{T}^{ODE}$ is
$\mathcal{J}_{0}(\theta^{*}_{\eta})$ $=-\big(\sigma_{Y_{T}^{ODE}}^{2}+(\mu_{Y_{T}^{ODE}}-1)^{2}\big)$
$=(-1)+(1+T^{2})^{-1}-\Bigg(1-(1+\frac{\beta}{2})^{-1}\cdot\frac{(1+T^{2})^{-\frac{1}{2}}-1}{(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-1}\Bigg)^{2}.$
Similarly, reward $\mathcal{J}_{SDE}$ for $Y_{T}^{SDE}$ is
$\mathcal{J}_{\eta}(\theta^{*}_{\eta})$ $=-\big(\sigma_{Y_{T}^{SDE}}^{2}+(\mu_{Y_{T}^{SDE}}-1)^{2}\big)$
$=(-1)+(1+T^{2})^{-(1+\eta^{2})}-\Big(1-(1+\frac{\beta}{2})^{-1}\Big)^{2}.$
For simplification, we denote
$\bar{T}:=1+T^{2}\in(T^{2},2T^{2})\,,\quad\bar{\beta}:=(1+\frac{\beta}{2})^{-1}\in(0,1].$
Now the reward gap
$\Delta_{\eta}$ $=\mathcal{J}_{SDE}-\mathcal{J}_{ODE}$
$=\Big((1+T^{2})^{-(1+\eta^{2})}-(1+T^{2})^{-1}\Big)$
$-\Bigg(\Big(1-(1+\frac{\beta}{2})^{-1}\Big)^{2}-\left(1-(1+\frac{\beta}{2})^{-1}\cdot\frac{(1+T^{2})^{-\frac{1}{2}}-1}{(1+T^{2})^{-\frac{1+\eta^{2}}{2}}-1}\right)^{2}\Bigg)$
$=\left(\frac{\bar{T}^{-\eta^{2}}-1}{\bar{T}}\right)-\left((1-\bar{\beta})^{2}-\left(1-\bar{\beta}\cdot\frac{\bar{T}^{-\frac{1}{2}}-1}{\bar{T}^{-\frac{1+\eta^{2}}{2}}-1}\right)^{2}\right).$
Since $0<\bar{T}^{-\eta^{2}}\leq 1$, we can bound
$\bar{T}^{-1}<\frac{\bar{T}^{-\eta^{2}}-1}{\bar{T}}\leq 0\,$,
and
$1-\bar{T}^{-\frac{1}{2}}=\frac{\bar{T}^{-\frac{1}{2}}-1}{0-1}<\frac{\bar{T}^{-\frac{1}{2}}-1}{\bar{T}^{-\frac{1+\eta^{2}}{2}}-1}\leq\frac{\bar{T}^{-\frac{1}{2}}-1}{\bar{T}^{-\frac{1}{2}}-1}=1.$
Therefore,
$|\Delta_{\eta}|$ $=\left|\frac{\bar{T}^{-\eta^{2}}-1}{\bar{T}}\right|+\left|(1-\bar{\beta})^{2}-\left(1-\bar{\beta}\cdot\frac{\bar{T}^{-\frac{1}{2}}-1}{\bar{T}^{-\frac{1+\eta^{2}}{2}}-1}\right)^{2}\right|$
$\leq\bar{T}^{-1}+\left|(1-\bar{\beta})^{2}-\left((1-\bar{\beta})+\bar{\beta}\cdot\bar{T}^{-\frac{1}{2}}\right)^{2}\right|$
$\leq\bar{T}^{-1}+\left(\bar{\beta}^{2}\cdot\bar{T}^{-1}+2\bar{\beta}(1-\bar{\beta})\bar{T}^{-\frac{1}{2}}\right)$
$\leq 2\bar{\beta}(1-\bar{\beta})T^{-1}+o(T^{-1})$
$\leq\frac{1}{2T}+o\left(\frac{1}{T}\right).$
Now we attempt to bound the reference gap. The reward $\mathcal{J}_{REF}$ for $Y_{T}^{REF}$ is
$\mathcal{J}_{\eta}(0)$ $=-\big(\sigma_{Y_{T}^{REF}}^{2}+(\mu_{Y_{T}^{REF}}-1)^{2}\big)$
$=(-1)+(1+T^{2})^{-(1+\eta^{2})}-1.$
Finally we have,
$\Delta_{\eta}^{REF}$ $=\mathcal{J}_{ODE}-\mathcal{J}_{REF}$
$=\bar{T}^{-1}-\left(1-\bar{\beta}\cdot\frac{\bar{T}^{-\frac{1}{2}}-1}{\bar{T}^{-\frac{1+\eta^{2}}{2}}-1}\right)^{2}-\left(\bar{T}^{-(1+\eta^{2})}-1\right)$
$\geq\bar{T}^{-1}-\left(2\bar{\beta}(1-\bar{\beta})\bar{T}^{-\frac{1}{2}}+o(\bar{T}^{-\frac{1}{2}})\right)-\left(\bar{T}^{-1}-1\right)$
$\geq 1-\frac{1}{2T}+o\left(\frac{1}{T}\right)$
C.2 Proof of Theorem 3.2
Similar to Appendix C.1, the quadratic reward (10) for $Y_{T}^{ODE}$ is
$\mathcal{J}_{ODE}$ $=-\big(\sigma_{Y_{T}^{ODE}}^{2}+(\mu_{Y_{T}^{ODE}}-1)^{2}\big)$
$=(-1)-\Bigg(1-(1+\frac{\beta}{2})^{-1}\cdot\frac{e^{-\frac{T^{2}}{2}}-1}{e^{-\frac{(1+\eta^{2})T^{2}}{2}}-1}\Bigg)^{2}.$
And the reward for $Y_{T}^{SDE}$ is
$\mathcal{J}_{SDE}$ $=-\big(\sigma_{Y_{T}^{SDE}}^{2}+(\mu_{Y_{T}^{SDE}}-1)^{2}\big)$
$=(-1)-\Big(1-(1+\frac{\beta}{2})^{-1}\Big)^{2}.$
With $\bar{\beta}:=(1+\frac{\beta}{2})^{-1}\in(0,1]$,
$|\Delta_{\eta}|$ $=\left|(1-\bar{\beta})^{2}-\left(1-\bar{\beta}\cdot\frac{e^{-\frac{T^{2}}{2}}-1}{e^{-\frac{(1+\eta^{2})T^{2}}{2}}-1}\right)^{2}\right|$
$\leq\left|(1-\bar{\beta})^{2}-\left((1-\bar{\beta})+\bar{\beta}\cdot e^{-\frac{T^{2}}{2}}\right)^{2}\right|$
$\leq\frac{e^{-\frac{T^{2}}{2}}}{2}+o\left(e^{-T^{2}}\right).$
Finally, $Y_{T}^{REF}\sim\mathcal{N}(0,1)$, so
$\mathcal{J}_{REF}=(-1)-(1-0)^{2}.$
And thus
$\Delta_{\eta}^{REF}$ $=\mathcal{J}_{ODE}-\mathcal{J}_{REF}$
$=1-\Bigg(1-(1+\frac{\beta}{2})^{-1}\cdot\frac{e^{-\frac{T^{2}}{2}}-1}{e^{-\frac{(1+\eta^{2})T^{2}}{2}}-1}\Bigg)^{2}$
$\geq 1-\frac{e^{-\frac{T^{2}}{2}}}{2}+o(e^{-T^{2}})$
C.3 Proof of Corollary 3.1
Let $\bullet\in\{SDE,ODE\}$. We decompose $Y^{\theta,\bullet}:=Y_{\parallel}^{\theta,\bullet}+Y_{\perp}^{\theta,\bullet}$ in terms of $\bm{r}$. Therefore,
$r(Y^{\theta,\bullet}(T))=-\left|Y_{\perp}^{\theta,\bullet}+(Y_{\parallel}^{\theta,\bullet}-\bm{r})\right|^{2}=-\left(\left|Y_{\perp}^{\theta,\bullet}\right|^{2}+\left|(Y_{\parallel}^{\theta,\bullet}-\bm{r})\right|^{2}\right)$
Since $\bm{\mu}_{i}\perp\bm{r}$, $\mathbb{E}[Y_{\parallel}^{\theta}(0)]=0$. Also, the backward dynamic injects a scaled multiple of $\bm{I}_{d}$ noise, so the coordinate-wise dynamics are independent. Therefore, we are able to analyze the $\text{Span}(\bm{r})$ subspace via separating each dimension $\bm{e}_{k}\in\text{Span}(\bm{r})$, in which $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},SDE})\right\}$ and $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},ODE})\right\}$ follows a similar controlled motion as discussed in Theorem 3.1 and 3.2. Moreover, the score function can be bounded by a constant of $\max_{i}{\sigma_{i}},\min_{i}^{-1}{\sigma_{i}}$ (Liang et al., 2025).
On the other hand, $\theta_{\perp}=\bm{0}$ is a minimizer for $\left|Y_{\perp}^{\theta,\bullet}\right|^{2}$, since an additional drift perpendicular to $\bm{r}$ does not alter reward variance but pushes away reward mean of $Y_{\perp}^{\theta,\bullet}(T)$ from reference priors. Therefore, $\theta=\theta_{\parallel}$. A similar analysis on Gaussian priors with controlled drifts for VP and VE dynamics yields to the desired bounds.
Appendix D Proof for Section 4
Proposition D.1.
(Gronwall’s) If $u^{\prime}(t)\leq\alpha(t)u(t)+\beta(t)$ and in addition satisfies the $\kappa$-strict negativity $\alpha(t)\leq-\kappa<0$ for some positive constant $\kappa$,
$u(T)\leq e^{-\kappa T}u(0)+\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds$ (17)
Proof.
The classical Gronwall’s theorem (Evans, 2010) gives
$u(T)\leq e^{\int_{0}^{T}\alpha(s)ds}\left(u(0)+\int_{0}^{T}e^{-\alpha(s)}\beta(s)ds\right).$
When $\alpha(t)\leq-\kappa$, $e^{\alpha(t)}\leq e^{-\kappa t}$ and the proposition follows naturally from integration. ∎
[…]
$u(T)\leq\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds.$ (18)
By Ito’s Lemma,
[…]
$=2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
$\qquad+g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
$\qquad-\eta^{2}g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{SDE})+\theta(t)\,\big\rangle+\eta^{2}g^{2}(t)$
$\leq 2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
$\qquad+||g||_{\infty}^{2}\cdot\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
[…]
By Condition 2 and 3,
[…]
Together with Young’s Inequality,
[…]
Since $\alpha(t)$ satisfies the $\kappa$-strict negativity by assumption, we may apply equation 18:
$u(T)$ $\leq\int_{0}^{T}e^{-\kappa(T-s)}\cdot\beta(s)ds$
[…]
Finally, $W_{2}$ distance is given by taking square root of $L^{2}$ distance between $Y^{ODE}$ and $Y^{SDE}$.
Appendix E Hyperparameters
All experiments are conducted on 7 Nvidia A100 GPUs. Mixed precision training is used with the bfloat16 (bf16) format.
E.1 DDPO Experiments
We follow the setup of Black et al. (2024), using denoising step $T=50$ and guidance weight $w=5.0$ throughout all experiments. We also use the AdamW optimizer Loshchilov & Hutter (2019) with default weight decay 1e-4 and optimal learning rates for different reward functions. Reward gaps under four reward functions are shown in Figure 3 in Section 5 and Appendix G.
Table 4: DDPO hyperparameters
ImageReward PickScore HPSv2 Aesthetic
DDPO Batch size (Per-GPU) 48 24 24 32
Samples per iteration (Global) 336 168 168 224
Gradient updates per iteration 2 2 2 4
Clip range 1e-5 5e-5 1e-4 1e-4
Optimizer Learning Rate 6e-4 6e-4 3e-4 3e-4
Animal Prompts dataset consists of 398 animal labels extracted from ImageNet-1k class labels Deng et al. (2009), often with comma-separated synonyms and scientific names.
Comprehensive Prompts dataset consists of 300 detailed and diverse descriptions of animals, vehicles, pieces of furniture, and landscapes with designated backgrounds, dynamics, or textiles.
E.2 MixGRPO Experiments
We follow the setup of Li et al. (2025), letting reward model be ”multi_\_reward” with equal weights. We set $T=15$ as the denoising steps, AdamW optimizer with learning rate 1e-5 and weight decay 1e-4. For GRPO, the generation group size is 12 and clip range is 1e-4. We perform 12 gradient updates per iteration.
Appendix F LLM Usage
Large Language Model (LLM) assists in LaTeX graphic alignments, spelling checks, and solving environment conflict issues in implementing DDPO and MixGRPO.
Appendix G More Experiment Results
G.1 DDPO Images on different training steps
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: SDE (top) and ODE (bottom) sampling from every 100 training steps under PickScore with $\eta=1.2$.
Prompts (from left to right): “African chameleon, Chamae”, “Gordon setter”, “Great Pyrenees”, “malamute, malemute, Alask, Bra”, “Siamese cat, Siamese”, “bee eater”, “gaint schnauzer”, “Indian elephant, Elephas”, “marmoset”, “water buffalo, water ox”, “pug, pug dog”.
G.2 DDPO Images under ImageReward
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Sampling schemes (from left to right): (i) SDE with $\eta=0.75$, (ii) ODE with $\eta=0.75$; (iii) SDE with $\eta=1.2$, (iv) ODE with $\eta=1.2$.
Prompts (from top to bottom): “collie”, “old English sheepdog, bob”, “Irish terrier”.
G.3 DDPO Images under PickScore
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Sampling schemes (from left to right): (i) SDE with $\eta=0.75$, (ii) ODE with $\eta=0.75$; (iii) SDE with $\eta=1.5$, (iv) ODE with $\eta=1.5$.
Prompts (from top to bottom): “baboon”, “white wolf, Arctic wolf”, “clumber, clumber spaniel”.
G.4 DDPO Reward Gaps for other rewards
G.4.1 ImageReward
Figure 9: Bounded reward gaps trained under ImageReward for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$
G.4.2 HPSv2
Figure 10: Bounded reward gaps trained under HPSv2 for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$
G.4.3 Aesthetic
Figure 11: Bounded reward gaps trained under Aesthetic for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$

### v2

@@
 Empirically, our findings through large-scale experiments on text-to-image models using denoising diffusion policy optimization (DDPO) and mixed group relative policy optimization (MixGRPO) validate that reward gaps consistently narrow over training, and ODE sampling quality improves when models are updated using higher-stochasticity SDE training.
+11footnotetext: Contacts: {js6646, hz2684, hc3136, yao, wt2319}@columbia.edu
 1 Introduction
-Diffusion models (e.g., Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), FLUX (Black Forest Labs, 2024)) have shown strong performance in text-to-image (T2I) tasks, and have also been extended beyond images to video (Ho et al., 2022) and audio (Liu et al., 2023).
-To meet downstream objectives such as aesthetics, safety, and alignment, it is essential to post-train with RLHF (Ouyang et al., 2022) for preference-driven improvements, often with a KL-regularization term to preserve performance on pretrained tasks (Schulman et al., 2017).
-Widely used RLHF algorithms include DDPO (Black et al., 2024) and GRPO (Shao et al., 2024) variants (FlowGRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), MixGRPO (Li et al., 2025)).
+Diffusion models (e.g., Stable Diffusion (stabledif), SDXL (sdxl), FLUX (flux)) have shown strong performance in text-to-image (T2I) tasks, and have also been extended beyond images to video (video) and audio (audio).
+To meet downstream objectives such as aesthetics, safety, and alignment, it is essential to post-train with RLHF (rlhf) for preference-driven improvements, often with a KL-regularization term to preserve performance on pretrained tasks (ppo).
+Widely used RLHF algorithms include DDPO (ddpo) and GRPO (grpo) variants (FlowGRPO (FlowGRPO), DanceGRPO (DanceGRPO), MixGRPO (MixGRPO)).
 DDPO directly optimizes human-preference rewards by casting the denoising process as a Markov Decision Process (MDP); GRPO variants use group-relative advantages.
-See (Winata et al., 2025) for a broader review of successful RLHF algorithms for generative models.
-Despite strong progress in alignment, RLHF training often exhibits unstable trajectories, long inference times, and vulnerability to reward hacking (Skalse et al., 2022).
-Using multiple rewards can mitigate the latter (Lee et al., 2025), but we also need efficient, robust samplers to produce stable, high-quality fine-tuned models.
-A classical scheme is DDPM (Ho et al., 2020), a discretization of the score-based backward SDE (Risken, 1996; Song et al., 2021b), which preserves data marginals but is time-consuming.
-In contrast, the deterministic DDIM sampler (Song et al., 2021a) follows the probability-flow ODE, enabling marginal-preserving sampling with fewer steps.
-DDPO typically uses DDIM with constant stochasticity to generate training samples; MixGRPO mixes SDE and ODE steps, varying stochasticity across the denoising horizon.
-However, although stochasticity is valuable for generating diverse training data, fine-tuned models often use DDIM or higher-order ODE solvers (Lu et al., 2022; 2025) for fast, stable inference.
-This naturally raises the question:
-Why can we guarantee good sampling quality when inference uses a different noise level than the one used during RLHF training?
-SDE Refer to caption Refer to caption Refer to caption Refer to caption
-ODE Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 1: ODE (below) image generation preserves prompt instructions with better quality on details compared to SDE (above) image generation under large stochasticity $(\eta=1.2)$.
-Prompts (from left to right): “A vintage writing desk with an open journal and a flickering candle.”, “A macaque soaking in a steaming hot spring, surrounded by falling snow.”, “A wicker rocking chair on a wrap-around porch during golden hour.”, “A sleek sports car drifting on a mountain highway during golden hour.”
-In this paper, we address it from three perspectives.
-Theoretically, we derive non-vacuous bounds on the reward gap between a generally SDE-fine-tuned model and its ODE-sampling counterpart using Gronwall’s inequality.
-Specifically, for VE and VP Gaussian models, the gap shrinks at sharp rates $O(1/T)$ and $O(e^{-T^{2}}/2)$, where $T$ is the denoising time horizon.
-These results justify the practice of ODE inference for SDE-fine-tuned models and provide insights into more randomized exploration beyond $\eta=1.0$, the noise level for continuous score-based backward dynamics.
-Methodologically, we adopt the gDDIM (Zhang et al., 2023) framework for arbitrary stochasticity levels.
-Rather than linear variance interpolations between DDIM and DDPM, our implementation preserves marginals for all $\eta\geq 0$ and, more importantly, supports $\eta>1$ in a principled way, enabling broader RLHF exploration beyond standard DDPM noise levels.
-Empirically, we evaluate large-scale T2I models and RLHF algorithms to quantify the reward gap across multiple reward functions.
-The results show that (i) reward gaps decrease as training quality improves, (ii) moderate-to-high training stochasticity (e.g., $\eta=1.2$) yields superior in- and out-of-domain performance, and (iii) ODE inference remains stable and often outperforms SDE inference under small denoising step budgets.
-  • Theory: Sharp bounds for VE/VP models; non-vacuous bounds for general distributions (linear in stochasticity, bounded in time-steps).
-  • Methodology: gDDIM-based noise scheduling that preserves marginals for all noise levels ($\eta\geq 0$) and supports high stochasticity ($\eta>1$) in a principled way.
-  • Experiments: Comprehensive evaluations across rewards and RLHF algorithms, validating bounded reward gaps and improved ODE inference with higher training stochasticity.
+RLHF training may also exhibit unstable trajectories, long inference times, and vulnerability to reward hacking (rewardhack; CMPO) and efficient, robust samplers (lu2022dpm; lu2025dpm) are desired for stable, high-quality fine-tuned models.
+See (winata2025preference) for a broader review of successful RLHF algorithms for generative models.
+The classical scheme DDPM (ddpm), a discretization of the score-based backward SDE (fp; Song20), allows for trajectory diversity and rich exploration in RLHF training; the deterministic DDIM sampler (ddim) follows a probability-flow ODE and facilitates inference.
+High stochasticity during training is crucial for effective exploration of the reward landscape.
+However, during inference, deterministic ODE sampling is preferred to ensure consistency and computational efficiency.
+This creates an inherent tension: we train via a highly stochastic process (SDE) but we deploy a deterministic process (ODE) for inference.
+This discrepancy necessitates the following question:
+To what extent can we guarantee ODE inference quality after we fine-tune diffusion models with SDE sampled trajectories?
+Figure 1: An illustration of the problem formulation.
+$\Delta_{\eta}$ represents the “Reward Gap” and $I_{\eta}$ represents the ODE reward improvement after SDE fine-tune.
+We would like $\Delta_{\eta}\downarrow$ with $I_{\eta}\uparrow$.
+Figure 1 illustrates the problem formulation.
+While SDE and ODE denoising processes (equation 3) yield the same terminal marginal for the pretrained model (see Section 2.1), a distributional difference arises between the distribution ($p_{\text{data}}^{\text{SDE}}$) after stochastic fine-tuning and the deterministic inference outcome ($p_{\text{data}}^{\text{ODE}}$) (see Section 3).
+We denote $\eta$ as the stochasticity scale, where $\eta=1.0$ corresponds to the standard variance in DDPM.
+We define this discrepancy as the Reward Gap ($\Delta_{\eta}$), which must be bounded to preserve the ODE reward improvement ($I_{\eta}$) gained over the baseline.
+In this paper, we present our analysis from three perspectives:
+  • Theory: We bound $\Delta_{\eta}$ between a generally SDE-fine-tuned model and its ODE-sampling counterpart using Gronwall’s inequality.
+Specifically, for VE, VP and mixture Gaussian models, the gap shrinks at sharp rates $O(1/T)$ and $O(e^{-T^{2}}/2)$, where $T$ is the denoising time horizon.
+  • Methodology: To handle high stochasticity beyond $\eta=1.0$ without changing marginals, we adopt the gDDIM scheme (gddim) for arbitrary stochasticity levels.
+This framework is introduced in Section 2.2 and 5.
+See Appendix E for more technical details.
+  • Experiments: We evaluate large-scale T2I models and RLHF algorithms, DDPO and MixGRPO, to quantify the reward gap across multiple reward functions.
+The results show that $\Delta_{\eta}$ decrease as training quality improves and $\eta=1.2$ yields superior ODE inference performance, $I_{\eta}$.
 Related Literature.
-Higher-order ODE solvers such as RX-DPM (Choi et al., 2025), DEIS (Zhang & Chen, 2022) approximate probability-flow.
-Recent analyses give broader convergence guarantees (Huang et al., 2025) and better representability (Chen et al., 2023) for score-based diffusion.
-In parallel, RL-based fine-tuning leverages stochastic sampling: DRaFT differentiates through noisy trajectories (Clark et al., 2023), Score-as-Action casts fine-tuning as stochastic control (Zhao et al., 2025), and Adjoint Matching enforces optimal memoryless noise schedules (Domingo-Enrich et al., 2024).
-Large-scale studies combine multiple rewards (Zhang et al., 2024); SEPO, D3PO, and ImageReFL boosts alignment (Zekri & Boullé, 2025; Yang et al., 2024; Sorokin et al., 2025).
-Finally, (Liang et al., 2025) gives discretization error bound and (Wu et al., 2024) discusses guidance (Ho & Salimans, 2021) for Gaussian Mixture priors.
+Higher-order ODE solvers such as RX-DPM (choi2025enhanced), DEIS (zhang2022fast) approximate probability-flow.
+Recent analyses give broader convergence guarantees (huang2025fast) and better representability (chen2023sampling) for score-based diffusion.
+In parallel, RL-based fine-tuning leverages stochastic sampling: DRaFT differentiates through noisy trajectories (clark2023directly), Score-as-Action casts fine-tuning as stochastic control (ZC25), and Adjoint Matching enforces optimal memoryless noise schedules (domingo2024adjoint).
+Large-scale studies combine multiple rewards (zhang2024large); SEPO, D3PO, and ImageReFL boosts alignment (zekri2025fine; yang2024using; sorokin2025imagerefl).
+Finally, (gaussianmixture) gives discretization error bound and (gm) discusses guidance (cfg) for Gaussian Mixture priors.
 Organization of the paper.
 The remainder of the paper is organized as follows.
-In Section 2, we provide background on diffusion models and the RLHF framework.
-The theory is developed in Sections 3 and 4.
-In Section 5, we conduct numerical experiments to justify the role of sampler stochasticity in fine-tuning large-scale T2I models.
-Concluding remarks are in Section 6.
+In Section 2, we provide background on diffusion RLHF and gDDIM in particular.
+Section 3 computes exact $\Delta_{\eta}$ and $I_{\eta}$ for simple controls.
+More complex networks are analyzed in Section 4.
+Numerical experiments in Section 5 validate our theoretical bounds by fine-tuning large-scale T2I models with DDPO and GRPO algorithms.
+Conclusion is given in Section 6.
 2 Preliminaries
-Figure 2: One-dimensional VP dynamics under different noise levels ($\eta$), using the same control function.
-Increasing $\eta$ introduces higher stochasticity in the trajectories while preserving the underlying dynamics.
 2.1 Forward-Backward Diffusion Models
-The continuous-time diffusion models (Song et al., 2021b) involve forward and backward procedures, and provide a framework that both unifies and extends earlier discrete-time models.
-The brief review below of this approach is based on (Tang & Zhao, 2025).
-The goal of diffusion models is to generate new samples that resemble the target data distribution $p_{\mbox{data}}(\cdot)$. The forward process is given by an SDE:
+The goal of diffusion models (Song20) is to generate new samples similar to $p_{\mbox{data}}(\cdot)$. The forward process is given by an SDE:
 $dX_{t}=f(t,X_{t})dt+g(t)dB_{t},\quad X_{0}\sim p_{\mbox{data}}(\cdot),$ (1)
-where $\{B_{t}\}$ is the $d$-dimensional Brownian motion, and $f:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are given functions (model parameters).
-Some regularity conditions on $f(\cdot,\cdot)$ and $g(\cdot)$ are required to ensure equation 1 is well-defined, see (Stroock & Varadhan, 1979).
-Moreover, assume the forward process $\{X_{t}\}$ has a smooth density $p(t,x):=\mathbb{P}(X_{t}\in dx)/dx$.
-It is known (Anderson, 1982; Haussmann & Pardoux, 1986) that under time reversal, the process, denoted $\overline{X}_{t}=X_{T-t}$, is governed by the SDE:
-$d\overline{X}_{t}=(-f(T-t,\overline{X}_{t})+g^{2}(T-t)\nabla\log p(T-t,\overline{X}_{t}))dt+g(T-t)dB_{t},\;\overline{X}_{0}\sim p(T,\cdot).$
-There are two key points to note.
-First, as the goal is to generate the target distribution, which is unknown, the backward process is initiated with a prior $p_{\mbox{noise}}(\cdot)$, a proxy of $p(T,\cdot)$), instead of setting $\overline{X}_{0}\sim p(T,\cdot)$.
-The choice of $p_{\mbox{noise}}(\cdot)$ depends on each specific model, and varies case by case in applications.
-Similarly, the choice of the pair of functions $(f,g)$ also depends on applications, and notable examples include variance exploding (VE) and variance preserving (VP) SDEs (see Section 3).
-Second, since the term $\nabla\log p(T-t,\overline{X}_{t})$ in the time-reversal process $\{\overline{X}_{t}\}$ is unknown, we need to find a way to estimate/approximate it.
-Specifically, calling $\nabla\log p(t,x)$ the “score function”, score matching refers to estimate the score function via a family of parameterized functions $\{s_{\theta}(t,x)\}_{\theta}$ (e.g., via neural nets).
-The most widely used approach is denoising score matching Vincent (2011):
-$\min_{\theta}\mathbb{E}_{t\sim\mbox{Unif}\,[0,T],\,X_{0}\sim p_{data}(\cdot)}[\mathbb{E}_{p(t,\cdot|X_{0})}|s_{\theta}(t,X_{t})-\nabla\log p(t,X_{t}|X_{0})|^{2}].$
-Replacing $\nabla\log p(t,x)$ by $s_{\theta}(t,x)$, the (backward) diffusion sampler is:
+where $\{B_{t}\}$ is the $d$-dimensional Brownian motion, and $f:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are given model parameters under regularity conditions (SV79).
+We use denoising score matching (Vi11) to minimize the mean square error between a parametrized neural network $s_{\theta}(t,X_{t})$ and the unknown term $\nabla\log p(T-t,\overline{X}_{t})$ in the time reversal SDE (HP86).
+The reverse-time SDE becomes:
 $dY_{t}=\left(-f(T-t,Y_{t})+g^{2}(T-t)s_{\theta}(T-t,Y_{t})\right)dt+g(T-t)dB_{t},\quad Y_{0}\sim p_{\mbox{noise}}(\cdot).$ (2)
-For richer levels of trajectory variation while keeping terminal marginals, the sampler is set to:
+For richer levels of sampler stochasticity, a flexible stochasticity parameter $\eta\geq 0$ is introduced:
 $dY_{t}=\left(-f(T-t,Y_{t})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta}(T-t,Y_{t})\right)dt+\eta\,g(T-t)dB_{t},\quad Y_{0}\sim p_{\mbox{noise}}(\cdot),$ (3)
-where $\eta\geq 0$ controls the sampler stochasticity along the trajectories.
-When $\eta=1$, equation 3 reduces to equation 2.
-When $\eta=0$, equation 3 becomes the probability flow ODE.
-We will use the processes specified by equation 3 as the pretrained models for fine-tuning.
-In particular, while prior studies focus on $\eta\in[0,1]$, we will explore the possible usage of $\eta>1$.
+With a divergence-free noise scale, equation 3 preserves the terminal marginals (Ander82).
 2.2 Discretization – Stochastic gDDIM
-As shown in (Song et al., 2021b), the DDPM scheme is a discretized SDE that preserves variance, and the deterministic DDIM is a discretization of the probability flow ODE that preserves variance.
-For most diffusion models, the model parameters are:
+As shown in (Song20), for most diffusion models, the model parameters are:
 $f(t,x)=\frac{1}{2}\frac{d\log\alpha_{t}}{dt}x\quad\mbox{and}\quad g(t)=\sqrt{-\frac{d\log\alpha_{t}}{dt}},$
-where $\{\alpha_{t}\}_{t=0}^{T}$ is a decreasing sequence with $\alpha_{0}=1$ and $\alpha_{T}=0$. (Zhang et al., 2023) proposed the following discretization of equation 3 for all $\eta\geq 0$:
+where $\{\alpha_{t}\}_{t=0}^{T}$ is a decreasing sequence from 1 to 0. For all $\eta\geq 0$, (gddim) proposed the generalized DDIM update:
 $x_{t-\Delta t}=\sqrt{\frac{\alpha_{t-\Delta t}}{\alpha_{t}}}x_{t}$ $+\left(\sqrt{\frac{\alpha_{t-\Delta t}}{\alpha_{t}}}(1-\alpha_{t})-\sqrt{(1-\alpha_{t-\Delta t}-\sigma_{t}^{2})(1-\alpha_{t})}\right)s_{\theta}(t,x_{t})$ (4)
@@
 $\sigma_{t}(\eta)=(1-\alpha_{t-\Delta t})\left(1-\left(\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}\right)^{\eta^{2}}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\right)^{\eta^{2}}\right).$ (5)
-equation 4 is referred to as the generalized denoising diffusion implicit model (gDDIM), which agrees with the stochastic DDIM (Song et al., 2021a).
-The difference lies in the choice of $\sigma_{t}(\eta)$: (Song et al., 2021a) relied on heuristics to suggest $\sigma_{t}(\eta)\propto\eta$, whereas equation 5 proposed by (Zhang et al., 2023) guarantees the exact sampling if the score function is known/precise.
-When $\eta=0$, we have $\sigma_{t}(0)=0$, and equation 4 is just the deterministic DDIM.
-We use gDDIM scheme to conduct numerical experiments with controlled stochasticity, of which the results are detailed in Section 5.
+For any $\eta>0$, the discrete scheme equation 4 has terminal marginal equal to equation 3.
+We use this discretization formulation for post-training with $\eta>1.0$ in Section 5.
+Details in Appendix E.
 2.3 Diffusion RLHF
-RLHF was originally proposed for LLM alignment (Bai et al., 2022; Ouyang et al., 2022).
-Such a paradigm can also be made well-suited for fine-tuning diffusion models, particularly to enhance T2I generation aligned with human feedback.
-The first works in this direction are (Black et al., 2024; Fan et al., 2023; Lee et al., 2023), which formulated the denoising step as Markov decision processes (MDPs).
-(Gao et al., 2024; Zhao et al., 2024; 2025) developed an RL approach based on continuous-time models.
-Here we focus on the most widely used approach of Denoising Diffusion Policy Optimization (DDPO).
-DDPO (Black et al., 2024): We formulate the denoising steps $\{x_{T},x_{T-1},\cdots,x_{0}\}$ as an MDP:
+RLHF was originally proposed for LLM alignment (Bai22; rlhf).
+(ddpo; dpok; Lee23) first formulated the denoising step as Markov decision processes (MDPs) to fine-tune diffusions.
+(Gao24; ZC24; ZC25) developed an RL approach based on continuous-time models.
+Here we focus on Denoising Diffusion Policy Optimization (DDPO) and Group Relative Policy Optimization (GRPO).
+DDPO (ddpo): We formulate the denoising steps $\{x_{T},x_{T-1},\cdots,x_{0}\}$ as an MDP:
 $s_{t}:=(\bm{c},t,x_{t}),$ $\pi(a_{t}|s_{t}):=p_{\theta}(x_{t-1}|x_{t},\bm{c}),\qquad$ $\mathbb{P}(s_{t+1}|s_{t},a_{t}):=\big(\delta_{\bm{c}},\delta_{t-1},\delta_{x_{t-1}}\big),$
 $a_{t}:=x_{t-1},$ $\rho_{0}(s_{0}):=\big(p(\bm{c}),\delta_{T},\mathcal{N}(0,I)\big),\qquad$ $R(s_{t},a_{t}):=\begin{cases}r(x_{0},\bm{c})&\text{if }t=0,\\ 0&\text{otherwise},\end{cases}$
-where $p_{\theta}$ is the probability from $x_{t}$ to $x_{t-1}$ given prompt $\bm{c}$ and $\theta$-parametrized score function $s_{\theta}$; $\delta_{\bullet}$ is the Dirac function that distributes the (non-zero) mass at $\bullet$ only; $R$ is the reward function that is equal to the denoised image quality $r(x_{0},c)$ for each trajectory.
-The objective of DDPO is to maximize:
+where $\delta_{\bullet}$ is the Dirac point mass concentrating at $\bullet$. The objective of DDPO is to maximize:
 $\mathcal{J}_{DDPO}(\theta):=\mathbb{E}_{\bm{c}\sim p(\bm{c}),x_{0}\sim p_{\theta}(x_{0}|\bm{c})}[r(x_{0},\bm{c})].$ (6)
-The policy gradient of equation 6 is $\nabla_{\theta}\mathcal{J}_{DDPO}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},\bm{c})\,r(x_{0},\bm{c})\right]$ (Williams, 1992).
-It is common to parameterize $p_{\theta}(x_{t-1}|x_{t},\bm{c})$ as isotopic Gaussian, so $\nabla_{\theta}\mathcal{J}_{DDPO}$ can be easily computed by Monte Carlo (Mohamed et al., 2020).
-GRPO (Shao et al., 2024): A recent breakthrough in the space of RLHF is the Group Relative Policy Optimization (GRPO) framework in the wake of Deepseek-R1 for reasoning.
-For T2I tasks, GRPO uses a group of sampled outputs (indexed by $i$ or $k$ below), along with the group normalization/average per prompt, to compute the advantages:
+The policy gradient of equation 6 is $\nabla_{\theta}\mathcal{J}_{DDPO}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},\bm{c})\,r(x_{0},\bm{c})\right]$ (reinforce).
+A common parameterization of $p_{\theta}$ is isotropic Gaussian (montecarlo).
+GRPO (grpo): GRPO first computes the advantages:
 $A_{i}:=\frac{r(x_{0}^{i},\bm{c})-\frac{1}{G}\sum_{k=1}^{G}r(x_{0}^{k},\bm{c})}{\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)},$
-where $G$ is the group size; $\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)$ denotes the standard deviation of the group rewards. The objective of GRPO is to maximize:
-$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=1}^{T}\,\min\left(\rho_{t}^{i}(\theta)A_{i},\text{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (7)
-where $\rho_{t}^{i}(\theta)=\frac{p_{\theta}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}{p_{\theta_{\mbox{old}}}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}$ is the likelihood ratio; $\{x_{t}^{i}\}$ is the $i$-th sample trajectory in the group; and the expectation is with respect to $\bm{c}\sim p(\bm{c})$ and $\{x^{i}_{t}\}\sim\pi_{\theta_{\mbox{old}}}(\cdot|\bm{c})$.
-Variants of the GRPO algorithm (Xue et al., 2025; Li et al., 2025) have been proposed for T2I tasks.
+where $G$ is the group size; $\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)$ denotes the standard deviation of group rewards. The objective of GRPO is to maximize:
+$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{N}\sum_{n=1}^{N}\,\min\left(\rho_{t}^{i}(\theta)A_{i},\text{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (7)
+where $\operatorname{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon):=\min\{\max\{\rho_{t}^{i}(\theta),1-\epsilon\},\,1+\epsilon\}$ restricts the likelihood ratio $\rho_{t}^{i}(\theta)=\frac{p_{\theta}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}{p_{\theta_{\mbox{old}}}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}$ within the range $1-\epsilon,1+\epsilon]$ by imposing hard constraints on the boundaries; $\{x_{t}^{i}\}$ is the $i$-th sample trajectory in the group; $N$ is the number of grouped outputs, and the expectation is with respect to $\bm{c}\sim p(\bm{c})$ and $\{x^{i}_{t}\}\sim\pi_{\theta_{\mbox{old}}}(\cdot|\bm{c})$.
 3 Theory on the reward gap
-As mentioned in the last section, the forward-backward diffusion approach to generative AI usually makes use of stochasticity for exploration (the forward step), while relying on deterministic samplers for inference (the backward step); the latter is obviously motivated by computational efficiency.
-Yet, there’s no study to examine/analyze the role of stochasticity; namely, whether there’s a right or “best” level of stochasticity to use, in both steps; and in particular, whether it’d be worthwhile to use stochasticity to the backward step, at the price of reduced efficiency.
-Denote by $\{Y^{REF}_{t}\}$ the “reference” model specified in equation 3), with a specific choice of the parameter $\eta$.
-Denote by $\{Y^{SDE}_{t}\}:=\{Y^{\theta}_{t}\}$ the fine-tuned model, score functions parameterized by $\theta$.
-To carry out this analysis, consider the following entropy-regularized reward objective (e.g., (Uehara et al., 2024; Tang, 2024)):
-$F_{\eta}(\theta)=\mathbb{E}[r(Y^{\theta}_{T})]-\beta\,\text{KL}(Y_{T}^{\theta}||Y_{T}^{REF}),$ (8)
-where $r(\cdot)$ is a reward function, and $\beta>0$ controls the level of exploration relative to the pretrained model.
-$Y^{SDE}_{t}$ evolves under a same $\eta$ as $Y^{REF}_{t}$, and the problem is to solve $\theta^{*}_{\eta}:=\arg\max_{\theta}F_{\eta}(\theta)$.
-$\bullet$ Inference: Denote by $\{Y_{t}^{ODE}\}$ the marginal-preserved (deterministic) ODE sampler by letting $\eta\downarrow 0$ in $\{Y^{SDE}_{t}\}$.
-Also let $\mathcal{J}_{\bullet}:=\mathbb{E}[r(Y_{T}^{\bullet})]$ be the evaluation metric, with $\bullet\in\{REF,SDE,ODE\}$.
-The goal is to understand how much improvement gained from fine-tuning, and the reward gap induced by the (deterministic) ODE sampler.
-This motivates the following definitions:
+As discussed in Section 1, forward–backward diffusion models typically employ stochastic SDE dynamics for exploration in the RLHF training, while adopting deterministic ODE samplers for inference (see Figure 1).
+In practice, we follow entropy-regularized objectives to fine-tune a parametrized family of score functions (UZ24; Tang24):
 Definition 3.1.
@@
   1. 1.
-    Improvement of fine-tuning as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{REF}_{t}\}$, $\mathcal{}I_{\eta}:=\mathcal{J}_{ODE}-\mathcal{J}_{REF}\geq 0$.
+    $\{Y_{T}^{\mathrm{REF}}\}$ as the samples extracted from the referenced base model following equation 3 with terminal marginal $p_{data}$;
   2. 2.
-    Reward gap as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{SDE}_{t}\}$, $\Delta_{\eta}:=|\mathcal{J}_{ODE}-\mathcal{J}_{SDE}|$.
-In what follows, we study $I_{\eta}$ and $\Delta_{\eta}$ for the VE/VP models with a Gaussian or mixture Gaussian target distribution, where the score function has a closed-form expression.
-A non-vacuous bound for the general case will be given in the next section.
+    We fine-tune the following objectives under a fixed stochasticity parameter $\eta$:
+    $F_{\eta}(\theta)=\mathbb{E}[r(Y_{T}^{\theta})]-\beta\,\mathrm{KL}(Y_{T}^{\theta}\,\|\,Y_{T}^{\mathrm{REF}}),\qquad\theta_{\eta}^{*}:=\arg\max_{\theta}F_{\eta}(\theta);$ (8)
+    let $\{Y_{t}^{\mathrm{SDE}}\}=\{Y_{t}(\theta_{\eta}^{*})\}$ be the optimal fine-tuned model, and $\{Y_{t}^{\mathrm{ODE}}\}$ be the deterministic sampler obtained by letting $\eta=0$:
+    $dY_{t}^{\text{SDE}}=\left(-f(T-t,Y_{t}^{\text{SDE}})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{\text{SDE}})\right)dt+\eta\,g(T-t)dB_{t},$ (9)
+    $dY_{t}^{\text{ODE}}=\left(-f(T-t,Y_{t}^{\text{ODE}})+\frac{1}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{\text{ODE}})\right)dt.$
+Although a pretrained model admits identical terminal marginals for arbitrary $\eta$ in the form of equation 3, once the data distribution is aligned to a human-preference reward $r(x)$, the distribution of $\{Y_{T}^{\mathrm{ODE}}\}$ and $\{Y_{T}^{\mathrm{SDE}}\}$ (we denote as $p_{data}^{\text{SDE}}$ and $p_{data}^{\text{ODE}}$) are not necessarily the same.
+Under a fixed noise level $\eta$, the terminal marginal satisfies a reward-tilting distribution $p_{data}^{\text{SDE}}\propto p_{\text{data}}(x)\exp(r(x)/\beta)$ (zhang2024rewardoveropt).
+The terminal score function
+$s_{\theta_{\eta}^{*}}(T,x)\approx\nabla\log p_{\theta_{\eta}^{*}}(T,x)=\nabla\log p_{\text{data}}(x)+\frac{\nabla r(x)}{\beta},$
+but the fine-tuned parameter $\theta_{\eta}^{*}$ does not regulate the intermediate trajectory, as the RLHF objective equation 8 only considers the terminal marginal.
+Therefore, we could not establish a general relationship between $s_{\theta_{\eta}^{*}}(t,x)$, $r(x)$, and $\nabla\log p_{\theta_{\eta}^{*}}(t,x)$ for $0<t<T$.
+In general,
+$\nabla\!\cdot\left(p_{\theta_{\eta}^{*}}(t,x)\cdot s_{\theta_{\eta}^{*}}(t,x)\right)\neq\Delta p_{\theta_{\eta}^{*}}(t,x).$
+As the reverse-time Fokker–Planck correspondence no longer satisfies the divergence free condition (Ander82), $p_{data}^{\text{SDE}}$ and $p_{data}^{\text{ODE}}$ need not coincide.
+Figure 2: One-dimensional VP dynamics under different noise levels $\eta$, using the same control function, demonstrating $\mathbb{E}(Y_{T}^{\text{SDE}})\approx\mathbb{E}(Y_{T}^{\text{ODE}})$ despite $p_{data}^{\text{SDE}}\neq p_{data}^{\text{ODE}}$.
+However, we are motivated by an observation in Figure 2 that 1D Variance Preserving (VP) dynamics, when run with the same constant downward drift, maintain a similar mean value regardless of the noise level $\eta$.
+we next rigorously quantify and generalize this observation Bounded reward gaps for simple dynamics are important for two reasons:
+  • Error Isolation.
+Under VP (as well as VE) models, we are able to compute the exact reward difference induced by changing $\eta$, free from the well-studied score-approximation and discretization error (gaussianmixture), with which we will also analyze in Section 4 and upper-bound reward gaps in those more complex settings.
+  • Parametrization Insight.
+Parametrization is inherent to the pre-trained (see equation 2) and the fine-tuned model (see equation 9) and affect RLHF performances in practice (han2025stochastic).
+Here we show that reward gap is small even under simple parametrization.
+As Section 4 bounds complex controls quadratic in $\eta$, our results are valid for general networks.
+Our goals are to quantify the reward improvement of fine-tuning $I_{\eta}$ and the reward gap $\Delta_{\eta}$ between SDE and ODE inference. Here we formally define:
+Definition 3.2.
+As illustrated in Figure 1, we define
+  1. 1.
+    Improvement of fine-tuning as the reward difference between $\{Y^{\text{ODE}}_{t}\}$ and $\{Y^{\text{REF}}_{t}\}$, $\mathcal{}I_{\eta}:=\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{REF}})]\geq 0.$
+  2. 2.
+    Reward gap as the reward difference between $\{Y^{\text{ODE}}_{t}\}$ and $\{Y^{\text{SDE}}_{t}\}$, $\Delta_{\eta}:=\left|\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{SDE}})]\right|.$
+In what follows, we analyze VE/VP models with a Gaussian or mixture Gaussian target distribution, where the score function has a closed-form expression without approximation and discretization errors, so the only discrepancy comes from changing $\eta$.
 Remark.
-$\mathcal{J}_{ODE}\geq\mathcal{J}_{REF}$ for all models, following from the optimality of $\theta^{*}_{\eta}$.
-$\mathcal{J}_{ODE}>\mathcal{J}_{SDE}$ for VE/VP models and DDPO; $\mathcal{J}_{SDE}>\mathcal{J}_{ODE}$ for MixGRPO.
+$\mathbb{E}[r(Y_{T}^{\text{ODE}})]\geq\mathbb{E}[r(Y_{T}^{\text{REF}})]$ holds for all models, following from the optimality of $\theta^{*}_{\eta}$.
+However, the sign of $\left(\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{SDE}})]\right)$ depends on the process dynamics as well as RLHF algorithms.
 3.1 VE with a Gaussian target
@@
 Since $X_{t}\sim\mathcal{N}(0,t^{2}+1)$, the (exact) score function is $\nabla\log p(t,x)=-\frac{x}{t^{2}+1}$. So the “pretrained” model is:
-$dY_{t}^{REF}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t}.$
-Next we set the reward function $r(x)=-(x-1)^{2}$, so the goal of fine-tuning is to drive the sample towards the mode $1$. We also specify the fine-tuned SDE and ODE by
-$dY_{t}^{SDE}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}(Y_{t}^{SDE}+\theta_{\eta}^{*})dt+\eta\sqrt{2(T-t)}dB_{t},$
-$dY_{t}^{ODE}=-\frac{T-t}{1+(T-t)^{2}}(Y_{t}^{ODE}+\theta_{\eta}^{*})dt,$
+$dY_{t}^{\text{REF}}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}Y_{t}^{\text{REF}}dt+\eta\sqrt{2(T-t)}dB_{t}.$
+Next we set the reward function $r(x)=-(x-1)^{2}$, so the goal of fine-tuning is to drive the sample towards mean $1$. We also specify the fine-tuned SDE and ODE as defined in equation 9 by
+$dY_{t}^{\text{SDE}}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}(Y_{t}^{\text{SDE}}+\theta_{\eta}^{*})dt+\eta\sqrt{2(T-t)}dB_{t},$
+$dY_{t}^{\text{ODE}}=-\frac{T-t}{1+(T-t)^{2}}(Y_{t}^{\text{ODE}}+\theta_{\eta}^{*})dt,$
 with $\theta_{\eta}^{*}$ maximizing the entropy-regularized reward (equation 8).
@@
 $r(x)=-(x-1)^{2}.$ (11)
-For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\right)\right]^{-1}$. Moreover,
+For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\right)\right]^{-1}$. Moreover, for $T\geq 1$
 $0\leq\Delta_{\eta}\leq\frac{1}{2T}+o\left(\frac{1}{T}\right)\quad\mbox{and}\quad\mathcal{I}_{\eta}\geq 1-\frac{1}{2T}+o\left(\frac{1}{T}\right).$ (12)
@@
 $dX_{t}=-tX_{t}dt+\sqrt{2t}\,dB_{t},\quad\mbox{with }\,p_{\mbox{data}}(\cdot)=\mathcal{N}(0,1).$ (13)
-Under the same setup as in the VE case, we have:
-$dY_{t}^{REF}=-\eta^{2}(T-t)Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t},$
-$dY_{t}^{SDE}=-\eta^{2}(T-t)Y_{t}^{SDE}\,dt-(1+\eta^{2})(T-t)\theta^{*}_{\eta}(t)\,dt+\eta\sqrt{2(T-t)}dB_{t},$
-$dY_{t}^{ODE}=-(T-t)\theta_{\eta}^{*}(t)dt,$
+Under the same setup as equation 9 and the VE case, we have:
+$dY_{t}^{\text{REF}}=-\eta^{2}(T-t)Y_{t}^{\text{REF}}dt+\eta\sqrt{2(T-t)}dB_{t},$
+$dY_{t}^{\text{SDE}}=-\eta^{2}(T-t)Y_{t}^{\text{SDE}}\,dt-(1+\eta^{2})(T-t)\theta^{*}_{\eta}(t)\,dt+\eta\sqrt{2(T-t)}dB_{t},$
+$dY_{t}^{\text{ODE}}=-(T-t)\theta_{\eta}^{*}(t)dt,$
 where a time-dependent control $\theta_{\eta}(t):=\theta_{\eta}e^{-\frac{(T-t)^{2}}{2}}$ is used for fine-tuning.
@@
 Theorem 3.2.
-Consider the Variance Preserving model (equation 12), with the reward function $r(x)=-(x-1)^{2}$.
+Consider the VP model (equation 13), with the reward function $r(x)=-(x-1)^{2}$.
 For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-e^{-\frac{(1+\eta^{2})T^{2}}{2}}\right)\right]^{-1}$.
-Moreover,
+Moreover, for $T\geq 1$
 $0\leq\Delta_{\eta}\leq\frac{e^{-T^{2}}}{2}+o\left(e^{-T^{2}}\right)\quad\mbox{and}\quad\mathcal{I}_{\eta}\geq 1-\frac{e^{-T^{2}}}{2}+o\left(e^{-T^{2}}\right).$ (14)
@@
 $\Delta_{\eta}\leq\left\{\begin{array}[]{ll}(2T)^{-1}&\text{for VE},\\ e^{-T^{2}}/2&\text{for VP}.\end{array}\right$. (15)
-For all these examples, the reward gap $\Delta_{\eta}\to 0$ (independent of $\eta$), as $T\to\infty$. Such a phenomenon will also be observed in fine-tuning T2I models with more complex rewards.
+For all these examples, the reward gap $\Delta_{\eta}\downarrow 0$ (independent of $\eta$), as $T\to\infty$.
+Such a phenomenon will also be observed in fine-tuning T2I models with more complex rewards.
 4 Non-vacuous bound on the reward gap
-In Section 3, we studied the reward gap for the VE and VP models with a Gaussian/mixture Gaussian target distribution.
-The analysis relies on the explicit computation of the score function, which is not available in general.
-Here our goal is to provide a bound on $W_{2}(Y_{T}^{SDE},Y_{T}^{ODE})$, where
-$dY_{t}^{SDE}=\left(-f(T-t,Y_{t}^{SDE})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{SDE})\right)dt+\eta\,g(T-t)dB_{t},$
-$dY_{t}^{ODE}=\left(-f(T-t,Y_{t}^{ODE})+\frac{1}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{ODE})\right)dt.$
-Here we fine-tune directly on the score matching function $\{s_{\theta}(t,x)\}_{\theta}$, instead of adding a control as in Section 3. We need the following assumptions.
+[…]
 Assumption 4.1.
-The following conditions hold:
+The following conditions hold for all $y,y_{1},y_{2}$ along inference trajectories,
   1. 1.
@@
 […]
-  3. 3.
-[…]
-The conditions 1 (dissipativity of $f$) and 2 (Lipschitz condition on $s_{\theta}$) are standard in stability analysis.
-The condition 3 ensures process $\{Y^{SDE}_{t}\}$ a contractive drift, which holds if $f$ is dissipative, and the fine-tuned distribution satisfies some strongly log-concave condition, see e.g., (Gao et al., 2025; Tang & Zhao, 2024).
-The following theorem gives a non‑vacuous bound on the reward gap, see proof in Appendix D.
+Condition 1 (Lipschitz of scores) and 2 ($L^{2}$ bound) are standard in stability analysis.
+Also, there always exists a continuous-time VP/VE SDE on $[0,1]$ whose discretization (ddim) matches the large-scale T2I models’ (i.e.
+Stable Diffusion and SDXL) training schedule.
+In addition, models (FLUX) based on rectified flow inherently uses a normalized time interval.
+Therefore, for large $\eta$, the following theorem yields $O(\eta^{2})$ bound on $W_{2}$ distance, see proof in Appendix D.1.
 $\eta$ ImageReward PickScore HPS_v2 Aesthetic
 Base
-0.320 20.69 0.253 5.375
+0.32 20.69 0.253 5.38
 ImageReward
-1.0 0.837 (0.918) 20.89  (20.94) 0.268 (0.271) 5.638  (5.720)
-1.2 0.915 (1.031) 20.95  (20.99) 0.268 (0.289) 5.697 (5.803)
-1.5 0.771 (0.907) 20.90 (20.95) 0.266 (0.269) 5.605 (5.735)
-PickScore 1.0 0.587 (0.729) 21.11 (21.28) 0.265 (0.267) 5.560 (5.678)
-1.2 0.594 (0.692) 21.17 (21.33) 0.264 (0.264) 5.568 (5.692)
-1.5 0.566 (0.701) 21.10 (21.36) 0.264 (0.262) 5.599 (5.769)
+1.0 0.84 (0.92) 20.89  (20.94) 0.268 (0.271) 5.64 (5.72)
+1.2 0.92 (1.03) 20.95  (20.99) 0.268 (0.289) 5.70 (5.80)
+1.5 0.77 (0.91) 20.90 (20.95) 0.266 (0.269) 5.61 (5.74)
+PickScore 1.0 0.59 (0.73) 21.11 (21.28) 0.265 (0.267) 5.56 (5.68)
+1.2 0.59 (0.69) 21.17 (21.33) 0.264 (0.264) 5.57 (5.69)
+1.5 0.57 (0.70) 21.10 (21.36) 0.264 (0.262) 5.60 (5.77)
 Table 1: Performance for ODE (SDE) samplers under DDPO fine-tuning. Bold numbers indicate the highest evaluations among all stochasticity, demonstrating $\eta=1.2$ in general performs the best.
 Theorem 4.1.
-[…]
-[…]
-[…]
-[…]
+If Assumption 4.1 hold with $T=1$, $W_{2}(Y_{T}^{\text{ODE}},Y_{T}^{\text{SDE}})\leq C\cdot\max\{\eta,\eta^{2}\}\cdot\sqrt{1+A}$, where $C$ only depends on $g||_{\infty}$ and $L$.
+Proposition 4.1.
+Let Assumption 4.1 hold with $L_{r}$-Lipschitz rewards $r$, $\Delta_{\eta}\leq L_{r}\cdot C\cdot\max\{\eta,\eta^{2}\}\cdot\sqrt{1+A}$, where $C$ only depends on $g||_{\infty}$ and $L$.
+An improved $O(\eta)$ bound relies on contractive coefficients (see Appendix D.2).
+This contractivity can arise from a contractive drift $f$ (TZ24).
+Alternatively, even though the backward dynamics of classical VP models are expansive, a strongly log-concave terminal distribution (GNZ23) still satisfies the contractive condition (see Assumption D.1), which can hold if the conditional terminal distribution for a fixed prompt $\bm{c}$ is approximately unimodal.
+Quadratic and linear growth rate on $\eta$ is consistent with our empirical observation: T2I fine-tuning preserves quality for ODE inference; rewards only deteriorate under very large $\eta$ (see Table 3).
+Remark.
+Discretization errors can be incorporated into the $W_{2}$ bound as an additional term $\epsilon_{d}(h)$ depending on the time–step size $h$ (gaussianmixture), with $\epsilon_{d}(h)\to 0$ as $h\to 0$.
+Under Assumption 4.1, Euler–Maruyama yields such a vanishing term; see Appendix D.3 for details.
 5 Numerical experiments
-In this section, we conduct numerical experiments for fine-tuning large-scale T2I models by leveraging sampler stochasticity, and examine the reward gap.
-Two classes of RL algorithms are considered: DDPO (Black et al., 2024) and MixGRPO (Li et al., 2025).
-The preference rewards that we use for fine-tuning DDPO include the LAION aesthetic (Schuhmann et al., 2022), HPS_v2.1 (Wu et al., 2023), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2023).
-The rewards for fine-tuning MixGRPO include HPS_v2.1, PickScore, ImageReward, and Unified Reward.
-(Wang et al., 2025) To study the reward gap, we use a high stochasticity $\eta$ under gDDIM scheme to generate training samples, and compare them with the ODE samples generated using the same weights in each iteration.
-The stochasticity scale $\eta$ controls the backward dynamics according to equation 4 and equation 5 in Section 2.2.
-Choices of hypermaraters are detailed in Appendix E.
+In this section, we fine-tune large-scale T2I models under large $\eta$ and examine its reward gap $\Delta_{\eta}$ with RLHF algorithms DDPO and MixGRPO (see Section 2.3).
+Preference rewards for fine-tuning DDPO include the LAION aesthetic (aesthetic), HPS_v2.1 (hpsv2), PickScore (pickscore), and ImageReward (imagereward).
+The rewards for fine-tuning MixGRPO include HPS CLIP, PickScore, ImageReward, and Unified Reward (unifiedreward).
+According to equation 4 and equation 5, we use a high stochasticity $\eta\geq 1.0$ under gDDIM scheme (see Section 2.2 and Appendix E) to generate robust training samples $\{Y_{T}^{\text{SDE}}\}$ and compare them with deterministic inference samples $\{Y_{T}^{\text{ODE}}\}$ following equation 9.
+The stochasticity scale $\eta$ controls the noise level of backward dynamics.
+Remark.
+In each experiment, the training stochasticity $\eta$ and the number of training steps $N$ are always specified. Choices of other hyperparameters are detailed in Appendix F.
 5.1 DDPO
-We use Stable Diffusion v1.5 (Rombach et al., 2022) as the base model, and fine-tune it with DDPO.
+We use Stable Diffusion v1.5 (stabledif) as the base model, and fine-tune it with DDPO.
 During training, we adopt ImageReward and PickScore as preference rewards, while Aesthetic and HPS_v2 are included as additional evaluation metrics.
-Figure 1 (and Appendix G) provide representative generations from the fine-tuned models.
+Figure 7 (and Appendix H) provides representative generations from the fine-tuned models.
 Our main observations are summarized as follows:
-$\bullet$ Decreasing Reward Gap with Quality Improvement: we conduct fine-tuning experiments under PickScore with $\eta=1.2$ and calculate SDE–ODE reward (contrary to Section 3, here we subtract ODE reward value from SDE reward for non-negativity) differences under multiple preference functions every 200 steps until reward collapse.
+$\bullet$ High Stochasticity Benefits Moderate Training Steps.
+We compare fine-tuning under ImageReward and PickScore at $\eta\in\{1.0,1.2,1.5\}$ for $N=200$ steps.
+As shown in Table 1, $\eta=1.2$ under ImageReward achieves the best in-domain and out-of-domain performance, while PickScore’s performances depend on evaluation metrics.
+$\bullet$ Decreasing Reward Gap with Quality Improvement.
+To study the reward gaps for smaller or larger $N$, we experiment under PickScore with $\eta=1.2$ and calculate SDE–ODE reward differences (here we report the SDE over ODE performance) under multiple preference functions every 200 steps until reward collapse.
 As shown in Table 2, the gap decreases as image quality improves for both samplers.
-$\bullet$ High Stochasticity Benefits Moderate Time Steps: we compare fine-tuning under ImageReward and PickScore at $\eta\in\{1.0,1.2,1.5\}$ for 200 steps.
-As shown in Table 1, $\eta=1.2$ under ImageReward achieves the best in-domain and out-of-domain performance, while PickScore’s performances depend on evaluation metrics.
-$\bullet$ Richer Prompt Contents Reduce Reward Gap: we compare performances with animal versus more comprehensive prompts (Figure 1, Appendix G) under ImageReward with $\eta=1.2$.
+$\bullet$ Richer Prompt Contents Reduce Reward Gap.
+We compare performances with animal versus more comprehensive prompts (see Appendix H.4) under ImageReward with $\eta=1.2$.
 As shown in Table 3, complex prompts generate higher post-tuning rewards and higher in-group variance in post-training.
 Moreover, their richer instructions reduce the SDE–ODE reward gap.
-Refer to caption Refer to caption
-Figure 3: Evolution of reward gap during DDPO training under PickScore fine-tuning with stochasticity $\eta\in\{1.0,1.2,1.5\}$.
-The gap for multiple results decreases steadily as training progresses, indicating that image quality improves for both samplers and that ODE inference remains competitive.
-$T=\bm{0}$ 200 400 600 800 1000 1200 1400 (1476)
+Figure 3: Evolution of reward gap (the reward difference between SDE sampling and ODE sampling) during DDPO training under PickScore fine-tuning with stochasticity $\eta\in\{1.0,1.2,1.5\}$.
+The gap for multiple rewards remains small as training step NN progresses.
+$N=\bm{0}$ 200 400 600 800 1000 1200 1400 (1476)
 ImgRwrd Gap 0.160 0.119 0.102 0.028 0.057 0.027 0.006 0.020 (0.564)
@@
 PickScore Gap 0.115 0.146 0.167 0.118 0.178 0.106 0.129 0.090 (0.907)
-SDE Reward 20.823 21.310 21.647 21.901 22.068 22.265 22.306 22.417 (18.920)
-ODE Reward 20.730 21.165 21.500 21.783 21.883 22.172 22.195 22.318 (17.044)
-Table 2: PickScore training until reward collapses. Smallest reward gaps and best sampler performances locate at the large training steps $T=1200,1400$.
+SDE Reward 20.82 21.31 21.65 21.90 22.07 22.27 22.31 22.42 (18.92)
+ODE Reward 20.73 21.17 21.50 21.78 21.88 22.17 22.20 22.32 (17.04)
+Table 2: PickScore training with $\eta=1.2$ until reward collapses. Smallest reward gaps and best sampler performances locate at the large training steps $N=1200,1400$.
 Animal Prompts Comprehensive Prompts
-$T=\bm{0}$ 100 200  $T=\bm{0}$ 100 200
-Mean 0.382 (0.539) 0.668 (0.805) 0.915 (1.031)    0.347 (0.470) 0.763 (0.872) 1.086 (1.174)
-Std 0.814 (0.799) 0.775 (0.721) 0.709 (0.652)    1.030 (0.997) 0.909 (0.868) 0.795 (0.727)
-Gap 0.146 0.138 0.116  0.115 0.110 0.106
-Table 3: Performance Comparison between prompts of different complexity. More complicated prompts yields faster fine-tuning improvements, larger in-group variances, and smaller SDE-ODE reward gaps.
+$N=\bm{0}$ 100 200  $N=\bm{0}$ 100 200
+Mean 0.38 (0.540) 0.67 (0.81) 0.92 (1.03)    0.35 (0.47) 0.76 (0.87) 1.09 (1.17)
+Std 0.81 (0.80) 0.78 (0.72) 0.71 (0.65)    1.03 (1.00) 0.91 (0.87) 0.80 (0.73)
+$\Delta_{\eta=1.2}$ 0.15 0.14 0.12  0.12 0.11 0.11
+Table 3: Performance Comparison between prompts of different complexity with $\eta=1.2$.
+More complicated prompts yields faster fine-tuning improvements, larger in-group variances, and smaller SDE-ODE reward gaps.
 5.2 MixGRPO
-We use FLUX.1 (Black Forest Labs, 2024) as the base model, and fine-tune it with MixGRPO, which is a sliding-window sampler that alternates between ODE and SDE schemes for 25 training steps in total.
+We use FLUX.1 (flux) as the base model, and fine-tune it with MixGRPO, which is a sliding-window sampler that alternates between ODE and SDE schemes for 25 training steps in total.
 Training is carried out with multiple rewards combined using equal weights, while evaluation is reported on ImageReward and HPSClip.
-$\bullet$ Bounded Reward Gap Under High Stochasticity: With $\eta=1.2$, ODE sampling in MixGRPO consistently outperforms the mixed SDE–ODE scheme, in contrast to the results from DDPO.
-As shown in the top-left and bottom-left panels of Figure 4, the reward gap steadily diminishes during training and converges to zero.
-$\bullet$ Quality Improvement: The middle and right panels of Figure 4 further shows that the ODE sampler performs consistently better on training prompts.
-For example, in the left pair of Figure 5, the generation with $\eta=1.2$ correctly aligns with the “trapped inside” prompt instruction, whereas the generation with $\eta=1.0$ fails to do so.
-Figure 4: Bounded reward gap (left column) and performance improvement (middle, right column) for MixGRPO
+$\bullet$ Bounded Reward Gap Under High Stochasticity.
+As shown in Figure 4, $\Delta_{\eta}$ converges to zero for both HPS Clip and ImageReward.
+Also, with $\eta=1.2$ and ImageReward metric, ODE sampling in MixGRPO consistently outperforms the mixed SDE–ODE scheme, in contrast to the results from DDPO.
+Figure 4: Bounded reward gap for MixGRPO, displayed with 7-step moving average. Here a positive score difference means the SDE component achieves higher reward than the ODE component.
+$\bullet$ T2I Quality Improvement.
+The upper (evaluated with ImageReward) and lower panels (evaluated with HPS Clip) of Figure 5 further show that samplers trained with larger stochasticity ($\eta=1.2$) perform consistently better than standard DDIM stochasticity ($\eta=1.0$) on training prompts.
+For example, in Figure 6, the generation with $\eta=1.2$ correctly aligns with the “trapped inside” prompt instruction, whereas the generation with $\eta=1.0$ fails to do so.
+Figure 5: Performance improvement for MixGRPO, displayed with 7-step moving average.
 Noise Level = 1.0 Refer to caption Refer to caption Refer to caption
@@
 6 Conclusion and Further Directions
-Sampling stochasticity in diffusion RLHF plays an important role yet remains underexamined.
-Here we present an explicit solution for generalized VE/VP diffusion models, together with extensive experiments, demonstrating that “high stochasticity in training samples, no stochasticity in generation” is both theoretically sound and practically advantageous.
-Our results open a broader space for tuning stochasticity hyperparameters, enabling more robust and diversified diffusion post-training.
-Future work will quantify the “reward gap” in video and multimodal generation and develop clearer theoretical insights into stochasticity as a drifting force toward the target distribution.
+This work clarifies the tension between stochastic SDE training and deterministic ODE inference in diffusion RLHF.
+By proving a bounded reward gap $\Delta_{\eta}$ and empirically showing that higher training stochasticity (e.g., $\eta=1.2$) improves deterministic image quality, we provide theoretical support for “training with SDE, inference with ODE”.
+Future work includes quantifying how distribution shift and the choice of reward function separately affect reward gaps.
+Acknowledgments
+Wenpin Tang and Jiayuan Sheng are supported by NSF grant DMS-2206038.
+Wenpin Tang acknowledges support by the Columbia Innovation Hub grant, and the Tang Family Assistant Professorship.
+The works of Haoxian Chen, Jiayuan Sheng, Hanyang Zhao, and David Yao are part of a ColumbiaCityU/HK collaborative project that is supported by InnoHK Initiative, The Government of the HKSAR and the AIFT Lab.
 Appendix A Lemmas on Linear Dynamics with Gaussian priors
@@
 Proof.
-A classical distance result on two Gaussian distributions (Hershey & Olsen, 2007) states:
+A classical distance result on two Gaussian distributions (4218101) states:
 Lemma B.1.
@@
 Therefore, we are able to analyze the $\text{Span}(\bm{r})$ subspace via separating each dimension $\bm{e}_{k}\in\text{Span}(\bm{r})$, in which $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},SDE})\right\}$ and $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},ODE})\right\}$ follows a similar controlled motion as discussed in Theorem 3.1 and 3.2.
-Moreover, the score function can be bounded by a constant of $\max_{i}{\sigma_{i}},\min_{i}^{-1}{\sigma_{i}}$ (Liang et al., 2025).
+Moreover, the score function can be bounded by a constant of $\max_{i}{\sigma_{i}},\min_{i}^{-1}{\sigma_{i}}$ (gaussianmixture).
 On the other hand, $\theta_{\perp}=\bm{0}$ is a minimizer for $\left|Y_{\perp}^{\theta,\bullet}\right|^{2}$, since an additional drift perpendicular to $\bm{r}$ does not alter reward variance but pushes away reward mean of $Y_{\perp}^{\theta,\bullet}(T)$ from reference priors.
@@
 Proposition D.1.
-(Gronwall’s) If $u^{\prime}(t)\leq\alpha(t)u(t)+\beta(t)$ and in addition satisfies the $\kappa$-strict negativity $\alpha(t)\leq-\kappa<0$ for some positive constant $\kappa$,
-$u(T)\leq e^{-\kappa T}u(0)+\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds$ (17)
-Proof.
-The classical Gronwall’s theorem (Evans, 2010) gives
-$u(T)\leq e^{\int_{0}^{T}\alpha(s)ds}\left(u(0)+\int_{0}^{T}e^{-\alpha(s)}\beta(s)ds\right).$
-When $\alpha(t)\leq-\kappa$, $e^{\alpha(t)}\leq e^{-\kappa t}$ and the proposition follows naturally from integration. ∎
+(Gronwall’s) Suppose integrable functions $u,\alpha,\beta:\mathbb{[}0,T]\to\mathbb{R}$ satisfies $u^{\prime}(t)\leq\alpha(t)u(t)+\beta(t)$ ,
+$u(T)\leq e^{\int_{0}^{T}\alpha(s)ds}\left(u(0)+\int_{0}^{T}e^{-\int_{0}^{s}\alpha(\tau)d\tau}\beta(s)ds\right).$ (17)
+A proof can be found in (evans).
+D.1 Proof of Theorem 4.1
 […]
-$u(T)\leq\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds.$ (18)
+$u(T)\leq e^{\alpha T}\left(\int_{0}^{T}\beta\cdot e^{-\alpha s}\,ds\right)=\frac{(e^{\alpha T}-1)\cdot\beta}{\alpha}.$ (18)
 By Ito’s Lemma,
 […]
-$=2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
-$\qquad+g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
-$\qquad-\eta^{2}g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{SDE})+\theta(t)\,\big\rangle+\eta^{2}g^{2}(t)$
-$\leq 2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
-$\qquad+||g||_{\infty}^{2}\cdot\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
-[…]
-By Condition 2 and 3,
-[…]
-Together with Young’s Inequality,
+$=2\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},-f(t,Y_{t}^{\text{ODE}})+f(t,Y_{t}^{SDE})\,\big\rangle$
+$\qquad\;+g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle$
+$\qquad\;-g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},\eta^{2}\cdot s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle+\eta^{2}g^{2}(t).$
+According to (TZ24tut),
+$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},-f(t,Y_{t}^{\text{ODE}})+f(t,Y_{t}^{\text{SDE}})\,\big\rangle=\left\{\begin{array}[]{ll}0&\text{for VE},\\ g^{2}(t)u(t)/2&\text{for VP}.\end{array}\right$.
+By Condition 1,
+$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle$ $\leq\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},L\cdot\left(Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}}\right))\,\big\rangle$
+$=L\cdot u(t).$
+By Condition 2,
+$\left|\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},\eta^{2}\cdot s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle\right|$ $\leq u(t)+\frac{\eta^{4}}{4}\cdot\mathbb{E}[||s_{\theta}(t,Y_{t}^{\text{SDE}})||^{2}$
+$=u(t)+\frac{\eta^{4}}{4}\cdot A.$
+Together, by absorbing $g||_{\infty}$ into an $\eta$-free constant $C_{\text{scheme}}$,
 […]
@@
 […]
+And we apply equation 18 with $T=1$:
+$u(T)$ $\leq\frac{(e^{\alpha T}-1)\cdot\beta}{\alpha}$
+$\leq\frac{\exp({C_{\text{scheme}}\cdot L})}{C_{\text{scheme}}\cdot L}\cdot\max\{\eta^{4},\eta^{2}\}\cdot\left(\frac{A}{4}+1\right).$
+Finally, the $W_{2}$ distance can be bounded in terms of this $L^{2}$–distance via the chosen coupling:
+$W_{2}\big(\mathcal{L}(Y_{T}^{\mathrm{ODE}}),\mathcal{L}(Y_{T}^{\mathrm{SDE}})\big)\;\leq\;\Big(\mathbb{E}\big[\|Y_{T}^{\mathrm{ODE}}-Y_{T}^{\mathrm{SDE}}\|^{2}\big]\Big)^{1/2}\;\leq\;C(\|g\|_{\infty},L)\,\max\{\eta,\eta^{2}\}\,\sqrt{1+A}.$
+D.2 Proof for strongly log-concave distributions
+Assumption D.1.
+(Strong log-concavity): Exists $\kappa>1$ such that for all $t,y_{1},y_{2}$:
 […]
-Since $\alpha(t)$ satisfies the $\kappa$-strict negativity by assumption, we may apply equation 18:
-$u(T)$ $\leq\int_{0}^{T}e^{-\kappa(T-s)}\cdot\beta(s)ds$
+Theorem D.1.
+[…]
+Proof.
+With the additional log-concavity assumption on the score function,
+$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle\leq-\kappa\cdot u(t).$
+In this case, the coefficient $\alpha$ becomes contractive as we may take $\delta=\frac{\kappa-1}{2\kappa}$:
 […]
 […]
-Finally, $W_{2}$ distance is given by taking square root of $L^{2}$ distance between $Y^{ODE}$ and $Y^{SDE}$.
-Appendix E Hyperparameters
+For arbitrary time horizon $T$, equation 18 gives:
+$u(T)$ $\leq\frac{(1-e^{-\alpha T})\cdot\beta}{-\alpha}$
+[…]
+[…]
+Remark.
+Assumption 4.2 is valid for a unimodal terminal distribution, in which a strongly log-concave coefficient $\kappa(T)$ satisfies
+$\mathbb{E}\big\langle\,Y_{T}^{\text{ODE}}-Y_{T}^{\text{SDE}},s_{\theta}(T,Y_{T}^{\text{ODE}})-s_{\theta}(T,Y_{T}^{\text{SDE}})\,\big\rangle\leq-\kappa(T)\cdot u(T).$
+In intermediate time steps,
+$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle\leq-\kappa(t)\cdot u(t),$
+in which (TZ24tut) bounds
+$\kappa(t)$ $\geq\frac{\kappa(T)}{\exp\left(-\int_{0}^{T-t}f(s)\,ds\right)+\kappa(T)\cdot\int_{0}^{T-t}\exp\left(-\int_{s}^{T-t}f(\tau)d\tau\right)ds}.$
+Since $\kappa(t)$ is strictly positive on $[0,T]$, a global coefficient $\kappa_{\text{global}}=\inf_{t}\kappa(t)$ exists.
+D.3 Discretization Error Analysis
+Let $N$ be the number of time steps and $h=T/N$ be the step size, we consider the DDPM sampling scheme with Euler–Maruyama discretization.
+Our Assumption 4.1 are equivalent to Assumptions 6.1 and 6.2 ($L$-Lipschitz score and finite second moment) in (gaussianmixture), which develops from (chen202).
+For $h\lesssim 1/L$,
+$\operatorname{TV}(q_{T}^{h},p_{0})\;\lesssim\;\underbrace{\sqrt{\mathrm{KL}(p_{0}\|\mathcal{N}(0,I))}e^{-T}}_{\text{forward convergence}}\;+\;\underbrace{(L\sqrt{dh}+Lm_{2}h)\sqrt{T}}_{\text{discretization error}}\;+\;\underbrace{\varepsilon_{0}\sqrt{T}}_{\text{score error}},$
+where $q_{T}^{h}$ is the law of the discrete sampler at time $T$, $d$ is the ambient space dimension, and $m_{2}^{2}=\mathbb{E}_{p_{0}}\|X\|^{2}$.
+In particular, the discretization contribution to the total variation distance is $O(h^{1/2})$ as $h\to 0$ for fixed $T,d,L,m_{2}$.
+Denote $Y_{T}^{\text{ODE},h}$ and $Y_{T}^{\text{SDE},h}$ as the ODE/SDE fine-tuned inference under step size $h$, by triangular inequality,
+$W_{2}(Y_{T}^{\text{ODE},h},Y_{T}^{\text{SDE},h})\leq W_{2}(Y_{T}^{\text{ODE}},Y_{T}^{\text{SDE}})+\underbrace{W_{2}(Y_{T}^{\text{ODE},h},Y_{T}^{\text{ODE}})+W_{2}(Y_{T}^{\text{SDE}},Y_{T}^{\text{SDE},h})}_{\epsilon_{d}(h)}$
+Since $\sup_{t}\mathbb{E}\|X_{t}\|^{2}\leq\infty$, a standard coupling method between $W_{2}$ and $\operatorname{TV}$ distance gives.
+$W_{2}(Y_{T}^{\theta},Y_{T}^{\theta,h})\;\;\lesssim\;\;\operatorname{TV}(Y_{T}^{\theta},Y_{T}^{\theta,h})^{1/2}.$
+Therefore, the contribution of time discretization to the $W_{2}$ error is bounded by the additive term
+$\epsilon_{d}(h):=C\,\bigl(L\sqrt{dh}+Lm_{2}h\bigr)^{1/2},$
+for some constant $C$ depending only on $T,d$. In the continuous time limit, $\epsilon_{d}(h)\to 0.$
+Appendix E Discussion on DDIM and gDDIM under high stochasticity
+When $1.0<\eta<+\infty$, the stochasticity level of gDDIM is always upper-bounded by
+$\sigma_{t}^{\text{gDDIM}}(\eta)=(1-\alpha_{t-\Delta t})\left(1-\left(\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}\right)^{\eta^{2}}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\right)^{\eta^{2}}\right)\leq 1-\alpha_{t-\Delta t}<+\infty;$
+whereas the linear interpolation of DDIM gives an unbounded
+$\sigma_{t}^{{\text{DDIM}}}(\eta)=\eta\sqrt{\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}.$
+In order to ensure the well-posedness of $\sqrt{(1-\alpha_{t-\Delta t}-\sigma_{t}^{2})(1-\alpha_{t})}$ in equation 4, the conventional discretization requires
+$\sigma_{t}^{\text{DDIM}}\leq 1-\alpha_{t-\Delta t}.$
+We may attempt to bypass the restriction by regulating
+$\sigma_{t}^{\text{linear DDIM}}(\eta)=\min\{1-\alpha_{t-\Delta t},\eta_{t}^{\text{DDIM}}(\eta)\}.$
+However, this changes the terminal marginals in equation 3.
+Therefore, the classical DDIM interpolation cannot simultaneously support arbitrary $\eta>1.0$ and preserve the terminal marginal, whereas gDDIM does.
+Appendix F Hyperparameters
 All experiments are conducted on 7 Nvidia A100 GPUs. Mixed precision training is used with the bfloat16 (bf16) format.
-E.1 DDPO Experiments
-We follow the setup of Black et al.
-(2024), using denoising step $T=50$ and guidance weight $w=5.0$ throughout all experiments.
-We also use the AdamW optimizer Loshchilov & Hutter (2019) with default weight decay 1e-4 and optimal learning rates for different reward functions.
-Reward gaps under four reward functions are shown in Figure 3 in Section 5 and Appendix G.
+F.1 DDPO Experiments
+We follow the setup of ddpo, using denoising step $T=50$ and guidance weight $w=5.0$ throughout all experiments.
+We also use the AdamW optimizer adamw with default weight decay 1e-4 and optimal learning rates for different reward functions.
+Reward gaps under four reward functions are shown in Figure 3 in Section 5 and Appendix H.
 Table 4: DDPO hyperparameters
@@
 Optimizer Learning Rate 6e-4 6e-4 3e-4 3e-4
-Animal Prompts dataset consists of 398 animal labels extracted from ImageNet-1k class labels Deng et al. (2009), often with comma-separated synonyms and scientific names.
+Animal Prompts dataset consists of 398 animal labels extracted from ImageNet-1k class labels imagenet, often with comma-separated synonyms and scientific names.
 Comprehensive Prompts dataset consists of 300 detailed and diverse descriptions of animals, vehicles, pieces of furniture, and landscapes with designated backgrounds, dynamics, or textiles.
-E.2 MixGRPO Experiments
-We follow the setup of Li et al.
-(2025), letting reward model be ”multi_\_reward” with equal weights.
+F.2 MixGRPO Experiments
+We follow the setup of MixGRPO, letting reward model be ”multi_\_reward” with equal weights.
 We set $T=15$ as the denoising steps, AdamW optimizer with learning rate 1e-5 and weight decay 1e-4.
@@
 We perform 12 gradient updates per iteration.
-Appendix F LLM Usage
+Appendix G LLM Usage
 Large Language Model (LLM) assists in LaTeX graphic alignments, spelling checks, and solving environment conflict issues in implementing DDPO and MixGRPO.
-Appendix G More Experiment Results
-G.1 DDPO Images on different training steps
+Appendix H More Experiment Results
+H.1 DDPO Images on different training steps
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Prompts (from left to right): “African chameleon, Chamae”, “Gordon setter”, “Great Pyrenees”, “malamute, malemute, Alask, Bra”, “Siamese cat, Siamese”, “bee eater”, “gaint schnauzer”, “Indian elephant, Elephas”, “marmoset”, “water buffalo, water ox”, “pug, pug dog”.
-G.2 DDPO Images under ImageReward
+H.2 DDPO Images under ImageReward
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Prompts (from top to bottom): “collie”, “old English sheepdog, bob”, “Irish terrier”.
-G.3 DDPO Images under PickScore
+H.3 DDPO Images under PickScore
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Prompts (from top to bottom): “baboon”, “white wolf, Arctic wolf”, “clumber, clumber spaniel”.
-G.4 DDPO Reward Gaps for other rewards
-G.4.1 ImageReward
+H.4 DDPO Images under more comprehensive prompts
+SDE Refer to caption Refer to caption Refer to caption Refer to caption
+ODE Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 10: ODE (below) image generation preserves prompt instructions with better quality on details compared to SDE (above) image generation under large stochasticity $(\eta=1.2)$.
+Prompts (from left to right): “A vintage writing desk with an open journal and a flickering candle.”, “A macaque soaking in a steaming hot spring, surrounded by falling snow.”, “A wicker rocking chair on a wrap-around porch during golden hour.”, “A sleek sports car drifting on a mountain highway during golden hour.”
+H.5 DDPO Reward Gaps for other rewards
+H.5.1 ImageReward
 Figure 11: Bounded reward gaps trained under ImageReward for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$
-G.4.2 HPSv2
+H.5.2 HPSv2
 Figure 12: Bounded reward gaps trained under HPSv2 for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$
-G.4.3 Aesthetic
+H.5.3 Aesthetic
 Figure 13: Bounded reward gaps trained under Aesthetic for 200 steps with stochasticity $\eta\in\{1.0,1.2,1.5\}$